### v2

# Moduli spaces of twisted equivariant $G$-bundles over a curve

Abstract.
Let $X$ be a compact Riemann surface, $\Gamma$ a finite group of automorphisms of $X$ and $G$ a connected reductive complex Lie group with center $Z$. If we equip this data with a homomorphism $\theta:\Gamma\to\operatorname{Aut}(G)$ and a 2-cocycle $c\in Z^{2}_{\theta}(\Gamma,Z)$, there is a notion of $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundle over $X$. The aim of this paper is to construct a coarse moduli space of isomorphism classes of polystable $(\theta,c)$-twisted equivariant $G$-bundles over $X$, according to the definition of polystability given by García-Prada–Gothen–Mundet i Riera. This generalizes the well-known construction of the moduli space of $G$-bundles given by Ramanathan. It also gives, in particular, a construction of the moduli space of $\Gamma$-equivariant $G$-bundles, and the moduli space of $G$-bundles for $G$ non-connected by our joint work with García-Prada–Gothen–Mundet i Riera — constructions which, to our knowledge, do not exist in the literature.
Key words and phrases:
Equivariant principal bundle, moduli spaces
2020 Mathematics Subject Classification:
Primary 14D20; Secondary 32L05
The author has received the support of the ERC Synergy Grant 810573/2019.
1. Introduction
Let $X$ be a compact Riemann surface, $\Gamma$ a finite group of automorphisms of $X$ and $G$ a connected reductive complex Lie group with center $Z$. Fix a homomorphism $\theta:\Gamma\to\operatorname{Aut}(G)$. Recall that there is a group of 2-cocycles $Z^{2}_{\theta}(\Gamma,Z)$, appearing for example in the context of Galois cohomology [23], consisting of maps $\Gamma\times\Gamma\to Z$ satisfying a precise condition (1). Pick a 2-cocycle $c\in Z^{2}_{\theta}(\Gamma,Z)$.
Given a $G$-bundle $Q\to X$, a $(\theta,c)$-twisted $\Gamma$-equivariant structure on $Q$ is a lift of the $\Gamma$-action on $X$ to a map from $\Gamma$ to the group of holomorphic automorphisms of the total space of $Q$. This action must further twist the bundle $G$-action via $\theta$, and the composition of operating by $\gamma_{1}$ and $\gamma_{2}\in\Gamma$ must be the same as multiplying by $c(\gamma_{1},\gamma_{2})$ and then operating by $\gamma_{1}\gamma_{2}$. The condition of $c$ being a 2-cocycle is equivalent to the associativity of this twisted equivariant structure.
A $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundle over $X$ is a pair $(Q,\bullet)$ consisting of a $G$-bundle $Q$ and a $(\theta,c)$-twisted $\Gamma$-equivariant structure $\bullet$ on $Q$. An isomorphism of twisted equivariants $G$-bundles is just a $\Gamma$-equivariant isomorphism of $G$-bundles.
A definition of $\zeta$-(poly, semi)-stability for twisted equivariant $G$-bundles — which are, in particular, twisted equivariant $G$-Higgs bundles — appears naturally in the work of García-Prada–Gothen–Mundet i Riera [13]. This definition fits into a Hitchin–Kobayashi correspondence for Higgs bundles with non-connected structure group, and is compatible with the description of fixed point subvarieties in moduli spaces principal bundles according to our joint work with García-Prada–Basu [2]. It depends on a parameter $\zeta$, which may be thought of as the topological type of the underlying $G$-bundle. This paper is dedicated to the construction of a projective variety parametrizing isomorphism classes of polystable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles, using Mumford’s GIT [19]. In particular, it shows that the (poly, semi)stability notions in the literature are also compatible with the corresponding GIT notions. The main result is the following.
Main Theorem (Theorem 7.7). There exists a complex projective variety $\mathcal{M}(X,G,\Gamma,\theta,c)$ parametrizing isomorphism classes of $\zeta$-polystable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$. It contains an open subvariety classifying isomorphism classes of stable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$.
In order to prove the main Theorem, we first construct a coarse moduli space for $G$ semisimple. In order to do this we need to formulate the problem in terms of GIT. First we fix an embedding $\iota:G\hookrightarrow\mathrm{GL}(n,\mathbb{C})$ such that $\Gamma$ acts on $G$ via inner automorphisms of $\mathrm{GL}(n,\mathbb{C})$. In Sections 3 and 4 we show that the category of (semi)stable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$ may be embedded into the category of $(\delta,\chi)$-(semi)stable pseudo-equivariant $\rho$-pairs, for $\chi\gg\delta\gg 0$ and a suitable representation $\rho:\mathrm{GL}(n,\mathbb{C})\to\mathrm{GL}(W)$. A pseudo-equivariant $\rho$-pair is a triple $(E,\sigma,f)$, where $E$ is a vector bundle of rank $n$ and trivial determinant, $\sigma:E(W)\to\mathcal{O}_{X}$ is a non-zero homomorphism and $f$ is a $\Gamma$-equivariant structure on $E$ satisfying certain compatibility condition (11) with $\sigma$. Here $E(W)$ is the associated vector bundle with fibre $W$. The notion of $(\delta,\chi)$-(semi)stability, which depends on two positive parameters $\delta$ and $\chi$, is defined using 1-parameter subgroups and weighted filtrations of $E$ — see Definition 4.4.
Using pseudo-equivariant pairs and Grothendieck’s Quot Scheme parametrizing quotients of $V\otimes\mathcal{O}_{X}$ of rank $n$ and trivial determinant for some fixed vector space $V$, we build a parameter space $\mathfrak{C}$ of pseudo-equivariant pairs containing all the (semi)stable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles. Moreover, $\mathfrak{C}$ is equipped with a $\mathrm{GL}(V)$-action such that two points are isomorphic if and only if they are in the same orbit — see Proposition 5.5. We then construct an injective morphism to a projective space $\mathbb{G}{(m)}$, called the Gieseker space — which depends on a positive integer $m$ — together with a compatible $\mathrm{GL}(V)$-action and a suitable linearization on some very ample line bundle — see Section 6.1. With these ingredients, Mumford’s GIT provides notions of (semi)stability which we show to be compatible with the notions of $(\delta,\chi)$-(semi)stability for pseudo-equivariant pairs if $\chi\gg\delta\gg 0$ and $m\gg 0$. We conclude that there exists a quasi-projective subvariety $\mathfrak{C}_{0}^{ss}\subset\mathfrak{C}$ of $\mathfrak{C}$ parametrizing semistable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles — see Corollary 6.7. The GIT quotient $\mathfrak{C}_{0}^{ss}\sslash\mathrm{GL}(V)$ is the required moduli space of polystable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles, for $G$ semisimple — see Theorem 6.11.
We then proceed to give a construction of a fine moduli space for $G$ abelian — see Theorem 7.6 —, and then use the semisimple and abelian cases to prove the main Theorem following a technique introduced by Ramanathan [22] and generalized by Gómez–Sols [14].
Here is the outline of the paper. In Section 2 we define $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles and the corresponding notions of $\zeta$-(poly, semi)stability. In Section 3 we prove an equivalence of categories between $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles and pseudo-equivariant $G$-bundles for $G$ semisimple. These are triples $(E,h,f)$ consisting of a vector bundle $E$ of rank $n$, a reduction of structure group $h$ of its bundle of frames to $G$ and a $\Gamma$-equivariant structure on $E$ satisfying a compatibility condition (8) with $h$. In Section 4.1 we define pseudo-equivariant $\rho$-pairs and the corresponding notion of $(\delta,\chi)$-(semi)stability. In Section 4.2 we embed the category of pseudo-equivariant $G$-bundles into the category of pseudo-equivariant $\rho$-pairs, where $\rho=\rho_{a,b,c}:\mathrm{GL}(n,\mathbb{C})\to\mathrm{GL}\left(\left(\left(\mathbb{C}^{n}\right)^{\otimes a}\right)^{\oplus b}\otimes(\bigwedge^{n}\mathbb{C}^{n})^{\otimes-c}\right)$ is induced by the standard representation, for some positive integers $a,b$ and $c$. In Section 4.3 we show that this restricts to an embedding of the category of (semi)stable pseudo-equivariant $G$-bundles into the category of $(\delta,\chi)$-(semi)stable pseudo-equivariant $\rho$-pairs, for $\chi\gg\delta\gg 0$ — see Proposition 4.11.
In Section 5 we build the parameter space $\mathfrak{C}$ of pseudo-equivariant pairs, equipped with a $\mathrm{GL}(V)$-action. In Sections 6.1 and 6.2 we find an injection into a Gieseker space, and a suitable linearization of the $\mathrm{GL}(V)$-action on a very ample line bundle such that GIT (semi)stability is equivalent to $(\delta,\chi)$-(semi)stability for $\chi\gg\delta\gg 0$. In Section 6.3 we show that the GIT quotient $\mathcal{M}(X,G,\Gamma,\theta,c):=\mathfrak{C}_{0}^{ss}\sslash\mathrm{GL}(V)$ is projective, where $\mathfrak{C}_{0}^{ss}\subset\mathfrak{C}$ is the subset of semistable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles. In Section 6.4 we show that closed $\mathrm{GL}(V)$-orbits in $\mathfrak{C}_{0}^{ss}$ are the orbits of polystable objects, hence $\mathcal{M}(X,G,\Gamma,\theta,c)$ parametrizes isomorphism classes of polystable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles.
In Section 7.2 we build a fine moduli space of polystable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles for $G$ abelian. This, together with the result for $G$ semisimple, allows us to construct the moduli space $\mathcal{M}(X,G^{\prime}\times Z,\Gamma,\theta,c)$, where $G^{\prime}$ is semisimple and $Z$ is abelian. Finally, using Lemma 7.3, in Section 7.3 we construct the moduli space $\mathcal{M}(X,G,\Gamma,\theta,c)$ of polystable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles for an arbitrary connected reductive complex Lie group $G$. The moduli space $\mathcal{M}(X,G,\Gamma,\theta,c)$ is given as a Galois cover of $\mathcal{M}(X,G^{\prime}\times Z,\Gamma,\theta,c)$, where $G^{\prime}=[G,G]$ and $Z$ is the centre of $G$.
Beyond its intrinsic geometric interest, we believe that a construction of the moduli space of polystable $(\theta,c)$-twisted equivariant $G$-bundles was overdue due to their important role in the literature. For example, the category of twisted equivariant $G$-bundles for $\theta$ and $c$ trivial is just the category of $\Gamma$-equivariant $G$-bundles. This has been extensively studied, for example in relation to parabolic bundles in the works of Balaji–Seshadri [1], Biswas [6] and Teleman–Woodward [24], and in the contex of the geometric Langlands programme in the works of Ben-Zvi–Francis–Nadler [5] and Chriss–Ginzburg [9].
When $c$ is trivial, the moduli stack of $\theta$-twisted $\Gamma$-equivariant $G$-bundles appears in relation with parahoric bundles in the work of Damiolini [10] and Damiolini–Hong [11]. Pappas–Rapoport [20] consider a more general setup where $G$ is replaced by a — non-necessarily constant — group scheme $\mathcal{G}^{\prime}$ over $X$, in relation to tamely ramified bundles. The category of $\theta$-twisted $\Gamma$-equivariant bundles also plays an important role in the study of the moduli space of $G$-Higgs bundles by Donagi–Gaitsgory [12].
In full generality, $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles appear naturally in the study of the moduli space of principal bundles with non-connected structure group and non-abelian Hodge theory — see our joint work with García-Prada–Gothen–Mundet i Riera [3], and the work of García-Prada–Gothen–Mundet i Riera [13] —, and also in the study of fixed point varieties of finite group actions on the moduli space of $G$-bundles according to our joint work with García-Prada–Basu [2].
Our results lay the groundwork for a deeper exploration of the geometry and topology of these moduli spaces, such as Harder–Narasimhan stratifications. An obvious complementary direction of study would be a stack-theoretic construction using beyond GIT techniques developed by Halpern–Leistner [15] and others. Another very relevant direction of research is the generalization to twisted equivariant Higgs pairs, which are very interesting objects from the point of view of non-abelian Hodge theory and the geometric Langlands programme as noted above. Finally, by allowing for the image of $\theta$ to contain anti-holomorphic automorphisms of $G$, one could enter the realm of pseudo-real $G$-bundles introduced by Biswas, García-Prada, Huismann and Hurtubise in [7] and [8].
2. Twisted equivariant $G$-bundles
Consider a compact Riemann surface $X$ and a connected reductive complex Lie group $G$ with center $Z$. Let $\Gamma$ be a finite group of automorphisms of $X$ acting on the right and take a homomorphism
$\theta:\Gamma\to\operatorname{Aut}(G);\,\gamma\mapsto\theta_{\gamma}.$
A 2-cocycle of $\Gamma$ with values in $Z$ is a map $c:\Gamma\times\Gamma\to Z$ which satisfies $c(\gamma,1)=c(1,\gamma)=1$ and
(1) $\theta_{\gamma_{0}}(c(\gamma_{1},\gamma_{2}))c(\gamma_{0},\gamma_{1}\gamma_{2})=c(\gamma_{0},\gamma_{1})c(\gamma_{0}\gamma_{1},\gamma_{2}).$
for every $\gamma,\gamma_{0},\gamma_{1},\gamma_{2}$ and $\gamma_{3}\in\Gamma$. Note that $\operatorname{Int}(G)$ acts trivially on $Z$, so (1) only depends on the outer class of $\theta$. The set of 2-cocycles forms a group which we denote by $Z^{2}_{\theta}(\Gamma,Z)$. There is an action of the group of maps $s:\Gamma\to Z$ on $Z^{2}_{\theta}(\Gamma,Z)$, namely
$Z^{\Gamma}\times Z^{2}_{\theta}(\Gamma,Z)\to Z^{2}_{\theta}(\Gamma,Z);\,(c,s)\mapsto s\cdot c:\Gamma\times\Gamma\to Z,$
$s\cdot c(\gamma_{1},\gamma_{2})=c(\gamma_{1},\gamma_{2})s(\gamma_{1})\theta_{\gamma_{1}}(s(\gamma_{2}))s(\gamma_{1}\gamma_{2})^{-1}.$
The quotient of $Z^{2}_{\theta}(\Gamma,Z)$ by this $Z^{\Gamma}$-action is the second cohomology group $H^{2}_{\theta}(\Gamma,Z)$. Fix a 2-cocycle $c\in Z^{2}_{\theta}(\Gamma,Z)$.
Given a holomorphic $G$-bundle $Q$ over $X$, a right $(\theta,c)$-twisted $\Gamma$-equivariant action on $Q$ is a holomorphic map
(2) $Q\times\Gamma\to Q;\,(q,\gamma)\mapsto p\bullet\gamma$
such that the following hold.
  • For every $q\in Q$, $q\bullet 1=q$.
  • For every $q\in Q$, $g\in G$ and $\gamma$ and $\gamma^{\prime}\in\Gamma$,
    (3) $(Qg)\bullet\gamma=(q\bullet\gamma)\theta_{\gamma}^{-1}(g)\quad\text{and}\quad(q\bullet\gamma)\bullet\gamma^{\prime}=(qc(\gamma,\gamma^{\prime}))\bullet(\gamma\gamma^{\prime}).$
  • The diagramme
    ${Q}$${Q}$${X}$${X}$$\scriptstyle{\bullet\gamma}$$\scriptstyle{\cdot\gamma}$
    commutes for each $\gamma\in\Gamma$.
Definition 2.1.
A $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundle over $X$ is a pair $(Q,\bullet)$ consisting of a holomorphic $G$-bundle $Q$ over $X$ and a $(\theta,c)$-twisted $\Gamma$-equivariant action $\bullet$ on $Q$. Whenever there is no danger of confusion we may omit $\theta,c,\Gamma$ or $\bullet$ from the notation. An isomorphism of $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles is a $\Gamma$-equivariant isomorphism of $G$-bundles.
Remark 2.2.
By [3, Proposition 2.16], the categories of $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles and $(\theta^{\prime},c^{\prime})$-twisted $\Gamma$-equivariant $G$-bundles are equivalent whenever $\theta$ and $\theta^{\prime}$ are in the same outer class, for some 2-cocycle $c^{\prime}$ depending on $c$ and $\theta$. The 2-cocycle $c^{\prime}$ is not unique, in fact we may set $\theta^{\prime}=\theta$ and choose $c^{\prime}$ to be any 2-cocycle cohomologous to $c$.
2.1. Semistable twisted equivariant $G$-bundles
[…]
(4) $P_{s}:=\{g\in G\;:\;e^{ts}ge^{-ts}\,\text{remains bounded as}\;t\to\infty\}.$
If $L_{s}$ is its Levi subgroup, then $K_{s}:=K\cap L_{s}$ is a maximal compact subgroup of $L_{s}$ and the embedding $K_{s}\hookrightarrow P_{s}$ is a homotopy equivalence. Now let $Q$ be a $G$-bundle with a holomorphic reduction $\tau\in H^{0}(X,Q(G/P_{s}))$, where $Q(G/P_{s})$ is the $G/P_{s}$-bundle associated to $E$ via the natural $G$-action; we denote by $Q_{\tau}:=\tau^{*}Q$ the corresponding $P_{s}$-bundle over $X$. Then there is a smooth reduction of structure group $\tau^{\prime}\in\Omega^{0}(X,Q_{\tau}/K_{s})$, and we may equip the corresponding $K_{s}$-bundle with a connection $A$ with curvature $F_{A}$. We define
(5) $\deg E(\tau,s):=\frac{i}{2\pi}\int_{X}\langle s,F_{A}\rangle.$
Note that, if $s$ is in the $\Gamma$-invariant part $i\mathfrak{k}^{\Gamma}$ of $i\mathfrak{k}$, then $P_{s}$ is $\Gamma$-invariant and so is $L_{s}$, since
(6) $L_{s}=\{g\in G\;:\;\lim_{t\to\infty}e^{ts}ge^{-ts}=g\}.$
In particular, $\theta$ induces $\Gamma$-actions on $G/P_{s}$ and $P_{s}/L_{s}$. If $Q$ is a $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundle then, combining the twisted equivariant action on $Q$ with the $\Gamma$-action on $G/P_{s}$, we obtain a $\Gamma$-equivariant structure on the fibre bundle $Q(G/P_{s})$ by [3, Proposition 3.16]. This lets us introduce the corresponding space of $\Gamma$-invariant reductions $H^{0}(X,Q(G/P_{s}))^{\Gamma}$. Given such a reduction $\tau\in H^{0}(X,Q(G/P_{s}))^{\Gamma}$, the corresponding $P_{s}$-bundle $Q_{\tau}$ inherits a $(\theta,c)$-twisted $\Gamma$-equivariant structure. Similarly, the fibre bundle $Q(P_{s}/L_{s})$ comes equipped with a $\Gamma$-equivariant action, with a corresponding space of $\Gamma$-invariant reductions $H^{0}(X,Q_{\tau}(P_{s}/L_{s}))^{\Gamma}$.
Definition 2.3.
Let $\mathfrak{z}_{\mathfrak{k}}$ be the centre of a $\Gamma$-invariant maximal compact subalgebra $\mathfrak{k}\subset\mathfrak{g}$, and fix $\zeta\in i\mathfrak{z}_{\mathfrak{k}}^{\Gamma}$. A $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundle $(Q,\bullet)$ over $X$ is:
  • $\zeta$-semistable if $\deg E(\tau,s)\geq\langle\zeta,s\rangle$ for any $s\in i\mathfrak{k}^{\Gamma}$ and any $\Gamma$-invariant reduction of structure group $\tau\in H^{0}(X,E(G/P_{s}))^{\Gamma}$.
  • $\zeta$-stable if $\deg E(\tau,s)>\langle\zeta,s\rangle$ for any $s\in i\mathfrak{k}^{\Gamma}$ and any reduction of structure group $\tau\in H^{0}(X,E(G/P_{s}))^{\Gamma}$.
  • $\zeta$-polystable if it is semistable and, for every $s\in i\mathfrak{k}^{\Gamma}$ and $\tau\in H^{0}(X,E(G/P_{s}))^{\Gamma}$ such that $\deg E(\tau,s)=\langle\zeta,s\rangle$, there is a further $\Gamma$-invariant holomorphic reduction of structure group $\tau^{\prime}\in H^{0}(X,E_{\tau}(P_{s}/L_{s}))^{\Gamma}$.
If $\zeta=0$, we omit it.
Remark 2.4.
If $\Gamma$ is the trivial group, we get the usual (poly, semi)stability notions for $G$-bundles.
In Section 5.1 we will need the following result, which is a special case of [4, Proposition 4.1].
Proposition 2.5.
A twisted equivariant $G$-bundle is $\zeta$-(poly)semistable if and only if the underlying $G$-bundle is $\zeta$-(poly)semistable.
3. Twisted equivariant $G$-bundles and pseudo-equivariant $G$-bundles
Pick a compact Riemann surface $X$, a finite group $\Gamma$ of complex automorphisms of $X$, a connected semisimple complex Lie group $G$ with center $Z$, a homomorphism $\theta:\Gamma\to\operatorname{Aut}(G)$ and a 2-cocycle $c\in Z^{2}_{\theta}(\Gamma,Z)$. The reason why we restrict to $G$ semisimple is that we first construct the moduli space under this assumption, and then extend our construction to arbitrary reductive $G$ in Section 7. In particular, since the centre $Z$ is discrete the stability parameter $\zeta=0$ and we omit it.
3.1. Pseudo-equivariant $G$-bundles
The twisted product $G\times_{(\theta,c)}\Gamma$ is the set $G\times\Gamma$ equipped with the group structure $(g,\gamma)(g^{\prime},\gamma^{\prime})=(g\theta_{\gamma}(g^{\prime})c(\gamma,\gamma^{\prime}),\gamma\gamma^{\prime})$ for every $(g,\gamma)$ and $(g^{\prime},\gamma^{\prime})$ in $G\times\Gamma$.
Since $G\times_{(\theta,c)}\Gamma$ is a finite extension of $G$ and $G$ is reductive, the group $G\times_{(\theta,c)}\Gamma$ is reductive. Therefore, there exists an embedding $\iota:G\times_{(\theta,c)}\Gamma\hookrightarrow\mathrm{GL}(n,\mathbb{C})$. Since $G$ is semisimple, $G=[G,G]$ and so $G$ has no non-trivial characters, which implies that $\iota(G)\subset\mathrm{SL}(n,\mathbb{C})$. We fix an embedding $\iota:G\times_{(\theta,c)}\Gamma\hookrightarrow\mathrm{SL}(n,\mathbb{C})$ once and for all. In particular every automorphism in $\theta(\Gamma)$ is realized as the restriction of an inner automorphism of $\mathrm{SL}(n,\mathbb{C})$, namely $\theta_{\gamma}=\operatorname{Int}_{(1,\gamma)}|_{G}$ for each $\gamma\in\Gamma$, where we write $(1,\gamma)$ instead of $\iota(1,\gamma)$ by abuse of notation. Here
$\operatorname{Int}_{(1,\gamma)}:\mathrm{GL}(n,\mathbb{C})\to\mathrm{GL}(n,\mathbb{C});\,g\mapsto(1,\gamma)g(1,\gamma)^{-1}$
denotes the conjugation automorphism.
Definition 3.1.
A $\Gamma$-pseudo-equivariant $G$-bundle of rank $n$ over $X$ is a triple $(E,h,f)$ consisting of a vector bundle $E$ of rank $n$, a reduction $h\in H^{0}(X,\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},E)/G)$ of structure group and a family of vector bundle isomorphisms $f=\{f_{\gamma}:E\xrightarrow{\sim}\gamma^{*}E\}_{\gamma\in\Gamma}$ satisfying
(7) $f_{1}=\operatorname{id},\;(\gamma^{*}f_{\gamma^{\prime}})\circ f_{\gamma}=f_{\gamma\gamma^{\prime}}$
and
(8) $f_{\gamma}(h)=(\gamma^{*}h)(1,\gamma)^{-1}.$
for every $\gamma$ and $\gamma^{\prime}\in\Gamma$. Here $\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},E)$ is the bundle of frames of $E$, and $f_{\gamma}(h)\in H^{0}(X,\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},\gamma^{*}E)/G)$ is the induced reduction of structure group. When there is no risk of confusion, we may omit $\Gamma$ from the notation. We sometimes call $f$ the pseudo-equivariant action or structure.
An isomorphism of pseudo-equivariant $G$-bundles $(E,h,f)\to(E^{\prime},h^{\prime},f^{\prime})$ is an isomorphism of vector bundles $\varphi:E\xrightarrow{\sim}E^{\prime}$ such that $\varphi(h)=h^{\prime}$ and $f_{\gamma}^{\prime}\circ\varphi=(\gamma^{*}\varphi)\circ f_{\gamma}$ for every $\gamma\in\Gamma$.
Remark 3.2.
For any pseudo-equivariant $G$-bundle $(E,h,f)$, the underlying vector bundle $E$ has trivial determinant, and in particular degree 00. This is due to the fact that the bundle of frames of $E$ has a reduction of structure group to $G\subset SL(n,\mathbb{C})$.
3.2. Twisted equivariant $G$-bundles as pseudo-equivariant $G$-bundles
Proposition 3.3.
The following categories are equivalent:
  1. (1)
    The category $\mathcal{A}$ of twisted equivariant $G$-bundles over $X$.
  2. (2)
    The category $\mathcal{B}$ of pseudo-equivariant $G$-bundles over $X$.
Proof.
We construct an invertible functor $\mathcal{A}\to\mathcal{B}$. Let $Q_{G}$ be a $G$-bundle equipped with a twisted equivariant $\Gamma$-action. By (3) this action is equivalent to a set of $G$-bundle isomorphisms $f_{\gamma}^{\prime}:Q_{G}\to\gamma^{*}\theta_{\gamma}(Q_{G})$ satisfying
(9) $f^{\prime}_{1}=1\quad\text{and}\quad(\gamma^{*}f^{\prime}_{\gamma^{\prime}})\circ f^{\prime}_{\gamma}=f^{\prime}_{\gamma\gamma^{\prime}}c(\gamma,\gamma^{\prime})$
for each $\gamma$ and $\gamma^{\prime}\in\Gamma$, where we identify elements of $\mathrm{GL}(n,\mathbb{C})$ with maps via the principal bundle action.
Let $Q:=Q_{G}(\mathrm{GL}(n,\mathbb{C}))$ be the extension of structure group and let $h\in H^{0}(X,Q/G)$ be the corresponding reduction of structure group. Recall that $\operatorname{Int}_{(1,\gamma)}(Q)$ may be defined as the $\mathrm{GL}(n,\mathbb{C})$-bundle that has the same total space as $Q$, but $\mathrm{GL}(n,\mathbb{C})$-action twisted by $\operatorname{Int}_{(1,\gamma)}^{-1}$. The set of isomorphisms $(f_{\gamma}^{\prime})_{\gamma\in\Gamma}$ induces a set of isomorphisms $f_{\gamma}:Q\to\gamma^{*}\operatorname{Int}_{(1,\gamma)}(Q)$ satisfying (9) and $f_{\gamma}(h)=\gamma^{*}h$ — here $h$ is regarded as a section of $\operatorname{Int}_{(1,\gamma)}(Q)$ via the identification of its underlying total space with $Q$.
Multiplication by $(1,\gamma)^{-1}$ — via the $\mathrm{GL}(n,\mathbb{C})$-action on $Q$ — provides an isomorphism $\gamma^{*}\operatorname{Int}_{(1,\gamma)}(Q)\to\gamma^{*}(Q)$, hence after composing we get a set of $\mathrm{GL}(n,\mathbb{C})$-bundle isomorphisms
$f_{\gamma}:=(1,\gamma)^{-1}\gamma^{*}f_{\gamma}:Q\xrightarrow{f_{\gamma}^{\prime}}\gamma^{*}\operatorname{Int}_{(1,\gamma)}(Q)\xrightarrow{(1,\gamma)^{-1}}\gamma^{*}(Q)$
sending $h$ to $\gamma^{*}h(1,\gamma)^{-1}$. Moreover, (9) implies (7) since, for every $q\in Q$,
$(\gamma^{*}f_{\gamma^{\prime}})\circ f_{\gamma}(q)$ $=(1,\gamma^{\prime})^{-1}\gamma^{*}f^{\prime}_{\gamma^{\prime}}\circ(1,\gamma)^{-1}f^{\prime}_{\gamma}(q)$
$=\gamma^{*}f^{\prime}_{\gamma^{\prime}}(f_{\gamma}^{\prime}(q)(1,\gamma)^{-1})(1,\gamma^{\prime})^{-1}$
$=\left((\gamma^{*}f^{\prime}_{\gamma^{\prime}})\circ f_{\gamma}^{\prime}(q)\right)\operatorname{Int}_{(1,\gamma^{\prime})}^{-1}((1,\gamma)^{-1})(1,\gamma^{\prime})^{-1}$
$=f^{\prime}_{\gamma\gamma^{\prime}}(qc(\gamma,\gamma^{\prime}))\left((1,\gamma^{\prime})\operatorname{Int}_{(1,\gamma^{\prime})}^{-1}((1,\gamma))\right)^{-1}$
$=f^{\prime}_{\gamma\gamma^{\prime}}(q)\operatorname{Int}_{(1,\gamma\gamma^{\prime})}^{-1}(c(\gamma,\gamma^{\prime}))\left((1,\gamma)(1,\gamma^{\prime})\right)^{-1}$
$=f^{\prime}_{\gamma\gamma^{\prime}}(q)(1,\gamma\gamma^{\prime})^{-1}c(\gamma,\gamma^{\prime})(1,\gamma\gamma^{\prime})\left((c(\gamma,\gamma^{\prime}),\gamma\gamma^{\prime})\right)^{-1}$
$=f^{\prime}_{\gamma\gamma^{\prime}}(q)(1,\gamma\gamma^{\prime})^{-1}c(\gamma,\gamma^{\prime})(1,\gamma\gamma^{\prime})(1,\gamma\gamma^{\prime})^{-1}c(\gamma,\gamma^{\prime})^{-1}$
$=f^{\prime}_{\gamma\gamma^{\prime}}(q)(1,\gamma\gamma^{\prime})^{-1}$
$=f_{\gamma\gamma^{\prime}}(q).$
Let $E:=Q(\mathbb{C}^{n})=Q\times_{\mathrm{GL}(n,\mathbb{C})}\mathbb{C}^{n}$ be the associated vector bundle of rank $n$. Then the isomorphisms $f_{\gamma}$ induce isomorphisms $f_{\gamma}:E\xrightarrow{\sim}\gamma^{*}E$, and $(E,h,f=\{f_{\gamma}\}_{\gamma\in\Gamma})$ is a pseudo-equivariant $G$-bundle. Therefore, starting with a twisted equivariant $G$-bundle $Q_{G}$, we have obtained a pseudo-equivariant $G$-bundle $(E,h,f)$. Furthermore, an isomorphism of twisted equivariant $G$-bundles induces, by extension of structure group, an isomorphism of pseudo-equivariant $G$-bundles. Working backwards we see that this functor is invertible. ∎
Definition 3.4.
A pseudo-equivariant $G$-bundle is called (semi)stable if the corresponding twisted equivariant $G$-bundle is so.
4. Pseudo-equivariant $G$-bundles as pseudo-equivariant $\rho$-pairs
Following the identification of $G$-bundles with swamps in [21, Section 2.4], we identify pseudo-equivariant $G$-bundles with analogous objects. We still assume that $G$ is a connected semisimple complex Lie group in this section. As in Section 3.1, fix the data $X$, $\Gamma$, $\theta$, $c$ and $\iota:G\times_{(\theta,c)}\Gamma\hookrightarrow\mathrm{SL}(n,\mathbb{C})$.
4.1. Semistable pseudo-equivariant pairs
Recall that a representation $\rho:\mathrm{GL}(n,\mathbb{C})\to\mathrm{GL}(W)$ is homogeneous if $\rho(z\operatorname{id})=z^{k}\operatorname{id}$ for some integer $k$, which is called the degree of $\rho$, or $k=\deg\rho$. Let $\rho:\mathrm{GL}(n,\mathbb{C})\to\mathrm{GL}(W)$ be a homogeneous representation. Assume that $W$ is also equipped with a right linear $\mathrm{GL}(n,\mathbb{C})$-action commuting with $\rho$, denoted by
(10) $W\times\mathrm{GL}(n,\mathbb{C})\to W;\,(w,g)\mapsto w\cdot g.$
Given a vector bundle $E\to X$ of rank $n$, denote by
$E(W):=\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},E)\times_{\rho}W$
the associated vector bundle with fibre $W$. Since the right $\mathrm{GL}(n,\mathbb{C})$-action (10) commutes with $\rho$, $E(W)$ is naturally equipped with a right $\mathrm{GL}(n,\mathbb{C})$-action.
Definition 4.1.
A $\Gamma$-pseudo-equivariant $\rho$-pair — abbreviated pseudo-equivariant pair when $\Gamma$ and $\rho$ are understood from context — is a triple $(E,\sigma,f)$ consisting of a vector bundle $E$ of rank $n$, a non-zero homomorphism $\sigma:E(W)\to\mathcal{O}_{X}$ and a family of vector bundle isomorphisms $f=\{f_{\gamma}:E\xrightarrow{\sim}\gamma^{*}E\}_{\gamma\in\Gamma}$ satisfying (7) — namely
$f_{1}=\operatorname{id}\quad\text{and}\quad(\gamma^{*}f_{\gamma^{\prime}})\circ f_{\gamma}=f_{\gamma\gamma^{\prime}}$
— and
(11) $f_{\gamma}(\sigma)=\gamma^{*}\sigma\cdot(1,\gamma)^{-1}.$
An isomorphism of pseudo-equivariant pairs $(E,\sigma,f)\xrightarrow{\sim}(E^{\prime},\sigma^{\prime},f^{\prime})$ consists of an isomorphism of vector bundles $\varphi:E\xrightarrow{\sim}E^{\prime}$ such that $(\gamma^{*}\varphi)\circ f_{\gamma}=f_{\gamma}^{\prime}\circ\varphi$ for each $\gamma\in\Gamma$ and $\varphi(\sigma)=\sigma^{\prime}$.
In order to define the stability notions for pseudo-equivariant $\rho$-pairs, we introduce some notation.
Definition 4.2.
Consider a vector space $V$ over a field $k$ and a representation $\widetilde{\rho}:\mathrm{GL}(n,k)\to\mathrm{GL}(V)$. A 1-parameter subgroup of $\mathrm{GL}(n,k)$ is a group homomorphism $\lambda:\mathbb{G}_{m}(k)\to\mathrm{GL}(n,k)$, where $\mathbb{G}_{m}(k)$ is the multiplicative group of $k$. Since $\mathbb{G}_{m}(k)$ is reductive, the representation $\widetilde{\rho}\circ\lambda$ is completely reducible, hence we can decompose $V$ as a sum of weight spaces $\bigoplus_{j=1}^{J}V^{\beta_{j}}$, where $t\in\mathbb{G}_{m}(k)$ acts on $V^{\beta_{j}}$ by multiplication by $t^{\beta_{j}}$ and $\beta_{j}\in\mathbb{Z}$. Given $v\in V$, we may decompose $v=\sum_{j=1}^{J}v^{\beta_{j}}$ uniquely, where $v^{\beta_{j}}\in V^{\beta_{j}}$, and define
(12) $\mu(v,\lambda):=-\min\{\beta_{j}\;:\;v^{\beta_{j}}\neq 0\}.$
Remark 4.3.
It is straightforward to check that, given two representations $\widetilde{\rho}_{i}:\mathrm{GL}(n,k)\to\mathrm{GL}(V_{i})$, two vectors $v_{i}\in V_{i}$ ($i=1,2$) and a 1-parameter subgroup $\lambda:\mathbb{G}_{m}(k)\to\mathrm{GL}(n,k)$,
$\mu(v_{1}\otimes v_{2},\lambda)=\mu(v_{1},\lambda)+\mu(v_{2},\lambda)\quad\text{and}\quad\mu(v_{1}\oplus v_{2},\lambda)=\max\left(\mu(v_{1},\lambda),\mu(v_{2},\lambda)\right).$
Fix a pseudo-equivariant pair $(E,\sigma,f)$ and let
$E_{\bullet}:\{0\}\subsetneq E_{1}\subsetneq\dots\subsetneq E_{J}\subsetneq E$
be a filtration of $E$ and let $\alpha_{\bullet}=(\alpha_{1},\dots,\alpha_{J})\in\mathbb{Q}_{>0}^{J}$. We call the pair $(E_{\bullet},\alpha_{\bullet})$ a weighted filtration of $E$, and we call it proper if $E_{\bullet}$ is proper. Then we set
(13) $M(E_{\bullet},\alpha_{\bullet}):=\sum_{j=1}^{J}\alpha_{j}(\deg(E)\operatorname{rk}(E_{j})-\deg(E_{j})\operatorname{rk}(E)).$
Set $\mathcal{H}om_{\Gamma}(E):=\bigoplus_{\gamma\in\Gamma}\mathcal{H}om(E,\gamma^{*}E)$. Let $\eta\in X$ be the generic point of $X$ when thought of as an algebraic curve. The restrictions $\mathbb{E}:=E|_{\eta}$, $\mathbb{W}:=E(W)|_{\eta}$ and $\mathbb{H}_{\Gamma}:=\mathcal{H}om_{\Gamma}(E)|_{\eta}$ are vector spaces over the function field $\mathbb{C}(X)$ of $X$. The restriction of the filtration $E_{\bullet}|_{\eta}$ is a filtration $\mathbb{E_{\bullet}}$ of $\mathbb{E}$ by vector subspaces. Let $r\in\mathbb{N}$ be such that $r\alpha_{\bullet}\in\frac{1}{n}\mathbb{Z}^{J}$. Let $\lambda:\mathbb{G}_{m}(\mathbb{C}(X))\to\mathrm{GL}(\mathbb{E})\cong\mathrm{GL}(n,\mathbb{C}(X))$ be a 1-parameter subgroup whose weight space decomposition is $\mathbb{E}=\bigoplus_{j=1}^{J}\mathbb{E}^{r\alpha_{j}}$ and satisfies $\mathbb{E}_{j}=\bigoplus_{i=1}^{j}\mathbb{E}^{r\alpha_{i}}$$\lambda$ is uniquely defined up to the conjugation action of the parabolic subgroup $\operatorname{Stab}_{\mathrm{GL}(\mathbb{E})}(\mathbb{E}_{\bullet})\leq\mathrm{GL}(\mathbb{E})$. Set $f^{*-1}:=\{f_{\gamma}^{*-1}:E^{*}\to\gamma^{*}E^{*}\}_{\gamma\in\Gamma}$,
(14) $\mu(E_{\bullet},\alpha_{\bullet},\sigma):=\frac{1}{r^{\deg\rho}}\mu(\sigma|_{\eta},\lambda)\quad\text{and}\quad\mu(E_{\bullet},\alpha_{\bullet},f):=\frac{1}{r}\max(\mu(\oplus f|_{\eta},\lambda),\mu(\oplus f^{*-1}|_{\eta},\lambda)),$
where $\sigma|_{\eta}\in\mathbb{W}$, and $\oplus f|_{\eta}$ and $\oplus f^{*-1}|_{\eta}\in\mathbb{H}_{\Gamma}$ — here we use the notation $\oplus f:=\bigoplus_{\gamma\in{\Gamma}}f_{\gamma}$. It is straightforward to check that (14) does not depend on $r$. By [21, Proposition 1.5.1.35] $\mu$ is invariant by the $\operatorname{Stab}_{\mathrm{GL}(n,\mathbb{C}(X))}(\mathbb{E}_{\bullet})\leq\mathrm{GL}(n,\mathbb{C}(X))$-action on $\lambda$, therefore (14) only depends on the data $(E_{\bullet},\alpha_{\bullet},\sigma,f)$.
Definition 4.4.
Given $\delta,\chi\in\mathbb{Q}_{>0}$, a pseudo-equivariant pair $(E,\sigma,f)$ is $(\delta,\chi)$-(semi)stable if, for every proper weighted filtration $(E_{\bullet},\alpha_{\bullet})$ of $E$,
(15) $M(E_{\bullet},\alpha_{\bullet})+\delta\mu(E_{\bullet},\alpha_{\bullet},\sigma)+\chi\mu(E_{\bullet},\alpha_{\bullet},f)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0.$
Remark 4.5.
Given a pseudo-equivariant $\rho$-pair $(E,\sigma,f)$, the underlying pair $(E,\sigma)$ defines a $\rho$-swamp $(E,\mathcal{O}_{X},\sigma)$ in the sense of [21]. Moreover, the swamp $(E,\mathcal{O}_{X},\sigma)$ is $\delta$-(semi)stable, in the sense of [21, Section 2.3.2], if and only if $(E,\sigma,f)$ is $(\delta,0)$-(semi)stable.
The following technical lemma will be useful when working with the notion of $(\delta,\chi)$-(semi)stability.
Lemma 4.6.
There exists a finite set
$S\subset\{(n_{\bullet},\alpha_{\bullet})\;:\;n_{j}\in\mathbb{N},\alpha_{j}\in\mathbb{Q}_{\geq 0},\sum_{j}n_{j}=n\},$
independent of $\delta$ and $\chi$, such that such that, for any pseudo-equivariant pair $(E,\sigma,f)$, the following two conditions are equivalent.
  1. (1)
    $(E,\sigma,f)$ is $(\delta,\chi)$-(semi)stable.
  2. (2)
    For any weighted filtration $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet}:=(\operatorname{rk}E_{j})_{j},\alpha_{\bullet})\in S$, (15) holds, namely
    $M(E_{\bullet},\alpha_{\bullet})+\delta\mu(E_{\bullet},\alpha_{\bullet},\sigma)+\chi\mu(E_{\bullet},\alpha_{\bullet},f)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0.$
Remark 4.7.
Lemma 4.6 may be summarized with the slogan $(\delta,\chi)$-semistability only has to be checked for weighted filtrations $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$.
Proof.
We follow the argument in [21, Example 1.5.1.18]. Choose a maximal torus $T\subset\mathrm{GL}(n,\mathbb{C}(X))$ with Lie algebra $\mathfrak{t}$ and a Weyl chamber $K\subset X(\mathfrak{t})$, where $X(\mathfrak{t})$ is the vector space of characters of $\mathfrak{t}$. Given a representation $\rho:\mathrm{GL}(n,\mathbb{C}(X))\to\mathrm{GL}(V)$, let $X(V)\subset X(\mathfrak{t})$ be the set of weights of $\rho$. Given a subset $R\subset X(V)$ and $\chi\in R$, set
$K(V,R,\chi):=\{k\in K\;:\;\langle k,\chi\rangle\leq\langle k,\chi^{\prime}\rangle\;\forall\chi^{\prime}\in R\}\subset X(\mathfrak{t}).$
This is a rational polyhedral cone in $\mathfrak{t}$, and
$K=\bigcup_{\chi\in R}K(V,R,\chi)$
is a fan decomposition of $K$.
Let
$V_{1}:=\mathbb{C}(X)^{n},\,V_{2}:=V\otimes\mathbb{C}(X)$
and
$V_{3}:=\bigoplus_{\gamma\in\Gamma}\operatorname{Hom}(\mathbb{C}(X),\gamma^{*}\mathbb{C}(X))\oplus\operatorname{Hom}(\mathbb{C}(X)^{*},\gamma^{*}\mathbb{C}(X)^{*}),$
equipped with the represenations $\mathrm{GL}(n,\mathbb{C}(X))\to\mathrm{GL}(V_{i})$ induced by the standard representation of $\mathrm{GL}(n,\mathbb{C}(X))$ — of course $\gamma^{*}\mathbb{C}(X)^{*}\cong\gamma^{*}\mathbb{C}(X)\cong\mathbb{C}(X)$ for every $\gamma$, but we are trying to use suggestive notation here. For every choice of subsets $R_{i}\subset X(V_{i})$, the expression
$K=\bigcup_{\chi_{i}\in R_{i}}\bigcap_{i=1,2,3}K(V_{i},R_{i},\chi_{i})$
is a fan decomposition of $K$. Let $\Lambda((V_{i})_{i=1,2,3},(R_{i})_{i=1,2,3},(\chi_{i})_{i=1,2,3})$ be the set of edges of $\bigcap_{i=1,2,3}K(V_{i},R_{i},\chi_{i})$, and let
$\Lambda:=\bigcup_{R_{i}\subset X(V_{i}),\chi_{i}\in R_{i}}\Lambda((V_{i}),(R_{i}),(\chi_{i})).$
Claim: Fix an isomorphism $X(\mathfrak{t})\cong X^{*}(\mathfrak{t})$ between the spaces of characters and cocharacters of $\mathfrak{t}$. Then, for every $v_{i}\in V_{i}$,
(16) $\mu(v_{1},\lambda)+\delta\mu(v_{2},\lambda)+\chi\mu(v_{3},\lambda)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$
for every non-trivial 1-parameter $\lambda$ of $\mathrm{GL}(n,\mathbb{C}(X))$, if and only if
(17) $\mu(g\cdot v_{1},\lambda)+\delta\mu(g\cdot v_{2},\lambda)+\chi\mu(g\cdot v_{3},\lambda)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$
for every $g\in\mathrm{GL}(n,\mathbb{C}(X))$ and every $\lambda$ corresponding to an element of $\Lambda$.
Since $\mu(g\cdot v_{i},\lambda)=\mu(v_{i},g^{-1}\lambda g)$, (16) implies 17. To prove the converse, assume (17) and pick $v_{i}\in V_{i}$. Let $\lambda:\mathbb{G}_{m}(\mathbb{C}(X))\to\mathrm{GL}(n,\mathbb{C}(X))$ be a non-trivial 1-parameter subgroup. There exists $g\in\mathrm{GL}(n,\mathbb{C}(X))$ such that $g\lambda g^{-1}$ lands in $T$. We may associate to $g\lambda g^{-1}$ a character $k$ which, after conjugating $\lambda$ further, we may assume to be in $K$. Let $R_{i}\subset X(V_{i})$ be the subset of weights appearing in the weight decomposition of $g\cdot v_{i}$. By (12),
$\mu(v_{i},\lambda)=\mu(g\cdot v_{i},g\lambda g^{-1})=-\langle k,\chi_{i}\rangle,$
where $\chi_{i}\in X(V_{i})$ satisfies that $\langle k,\chi_{i}\rangle\leq\langle k,\chi_{i}^{\prime}\rangle$ for every $\chi_{i}^{\prime}\in R_{i}$. In other words, $k\in K(V_{i},R_{i},\chi_{i})$ for $i=1,2,3$, or $k\in\bigcap_{i}K(V_{i},R_{i},\chi_{i})$. There is a decomposition
$k=\sum_{k^{\prime}\in\Lambda((V_{i})_{i},(R_{i})_{i},(\chi_{i})_{i})}l_{k^{\prime}}^{k}k^{\prime},$
for some positive numbers $l_{k^{\prime}}^{k}$. Then,
$\mu(v_{i},\lambda)=-\sum_{k^{\prime}\in\Lambda((V_{i})_{i},(R_{i})_{i},(\chi_{i})_{i})}l_{k^{\prime}}^{k}\langle k^{\prime},\chi_{i}\rangle.$
Note that, if we regard $k^{\prime}$ as a 1-parameter subgroup by abuse of notation, via the chosen isomorphism $X(\mathfrak{t})\cong X^{*}(\mathfrak{t})$, then $\mu(g\cdot v_{i},k^{\prime})=-\langle k^{\prime},\chi_{i}\rangle$, since $k^{\prime}\in K(V_{i},R_{i},\chi_{i})$. Therefore,
$\mu(v_{i},\lambda)=\sum_{k^{\prime}\in\Lambda((V_{i})_{i},(R_{i})_{i},(\chi_{i})_{i})}l_{k^{\prime}}^{k}\mu(g\cdot v_{i},k^{\prime}),$
and so
$\mu(v_{1},\lambda)+\delta\mu(v_{2},\lambda)+\chi\mu(v_{3},\lambda)=$
$\sum_{k^{\prime}\in\Lambda((V_{i})_{i},(R_{i})_{i},(\chi_{i})_{i})}l_{k^{\prime}}^{k}\left(\mu(g\cdot v_{1},k^{\prime})+\delta\mu(g\cdot v_{2},k^{\prime})+\chi\mu(g\cdot v_{3},k^{\prime})\right)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0,$
where the last inequality follows from $\Lambda((V_{i})_{i},(R_{i})_{i},(\chi_{i})_{i})\subset\Lambda$ and $l_{k^{\prime}}^{k}\geq 0$, as required.
To finish the proof, we identify $K$ with $\mathbb{R}^{n}_{\geq 0}$, with basis $(e_{1},\dots,e_{n})$. For each $k\in\Lambda$, let $\beta_{j}^{k}\geq 0$ be such that
$k=\sum_{j}\beta^{k}_{j}e_{j}.$
Consider the set of tuples $\{(\beta^{k}_{j})_{j}\}\subset\mathbb{R}^{n}_{\geq 0}$. There is a map
$\{(\beta^{k}_{j})_{j}\}\to\{(n_{\bullet},\alpha_{\bullet})\;:\;n_{j}\in\mathbb{N},\alpha_{j}\in\mathbb{Q}_{\geq 0},\sum_{j}n_{j}=n\},$
given by ordering each tuple $(\beta^{k}_{j})_{j}$, partitioning the tuple into sets of equal numbers and letting $n_{\bullet}$ being the data of the partition and $\alpha_{\bullet}$ represent the value of $(\beta^{k}_{j})$ on each block. Denoting the image by $S$, the lemma follows from the claim. ∎
4.2. Pseudo-equivariant $G$-bundles as pseudo-equivariant pairs
Let $(E,h,f)$ be a pseudo-equivariant $G$-bundle. The commutative diagramme
${\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},E)}$${\mathcal{H}om(\mathcal{O}_{X}^{\oplus n},E)}$${\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},E)/G}$${\mathcal{H}om(\mathcal{O}_{X}^{\oplus n},E)\sslash G,}$
where the double slash denotes GIT quotient, implies that the total space of $\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},E)/G$ is an open subvariety of the total space of $\mathcal{H}om(\mathcal{O}_{X}^{\oplus n},E)\sslash G$. Indeed, since $G$ is semisimple, $\iota(G)\subset\mathrm{SL}(n,\mathbb{C})$, hence the determinant morphism $\det:\mathcal{H}om(\mathcal{O}_{X}^{\oplus n},E)\to\mathcal{O}_{X}$ is $G$-equivariant and descends to a morphism
$\det:\mathcal{H}om(\mathcal{O}_{X}^{\oplus n},E)\sslash G\to\mathcal{O}_{X}.$
Moreover, $\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},E)/G=\det^{-1}(\mathcal{O}_{X}^{*})$, where $\mathcal{O}_{X}^{*}\subset\mathcal{O}_{X}$ denotes the subsheaf of functions with no zeros, and
(18) $H^{0}(X,\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},E)/G)=H^{0}(X,\mathcal{H}om(\mathcal{O}_{X}^{\oplus n},E)\sslash G)\setminus\det{}^{-1}(0),$
where 00 denotes the zero section of $\mathcal{O}_{X}$.
On the other hand $\mathcal{H}om(\mathcal{O}_{X}^{\oplus n},E)\sslash G=\mathcal{S}pec(\mathcal{S}ym^{*}(\mathcal{O}_{X}^{\oplus n}\otimes E^{*})^{G})$, where $\mathcal{S}ym^{*}(\mathcal{O}_{X}^{\oplus n}\otimes E^{*})^{G}\subset\mathcal{S}ym^{*}(\mathcal{O}_{X}^{\oplus n}\otimes E^{*})$ is the subsheaf of $G$-invariant sections, hence a global section of the sheaf $\mathcal{H}om(\mathcal{O}_{X}^{\oplus n},E)\sslash G$ is just a homomorphism of sheaves $\mathcal{S}ym^{*}(\mathcal{O}_{X}^{\oplus n}\otimes E^{*})^{G}\to\mathcal{O}_{X}$. Note that $\mathcal{S}ym^{*}(\mathcal{O}_{X}^{\oplus n}\otimes E^{*})^{G}$ is finitely generated over $\mathcal{O}_{X}$, therefore, by [18, Chapter III, Section 8, Lemma], there exists a positive integer $d$ such that $\mathcal{S}ym^{*d}(\mathcal{O}_{X}^{\oplus n}\otimes E^{*})^{G}$ is generated by its degree 1 part, that is $\mathcal{S}ym^{d}(\mathcal{O}_{X}^{\oplus n}\otimes E^{*})^{G}$. Via the Veronese embedding, $\mathcal{S}ym^{*}(\mathcal{O}_{X}^{\oplus n}\otimes E^{*})^{G}\cong\mathcal{S}ym^{*d}(\mathcal{O}_{X}^{\oplus n}\otimes E^{*})^{G}$. Therefore, a homomorphism $h:\mathcal{S}ym^{*}(\mathcal{O}_{X}^{\oplus n}\otimes E^{*})^{G}\to\mathcal{O}_{X}$ is completely determined by a homomorphism
(19) $h:\mathcal{S}ym^{d}(\mathcal{O}_{X}^{\oplus n}\otimes E^{*})^{G}\to\mathcal{O}_{X}$
which we also denote by $h$ by abuse of notation.
Next we want to express $h$ as a homomorphism $E(W)\to\mathcal{O}_{X}$, where $W$ satisfies the assumptions of Section 3.1. First we need a technical lemma. Given positive integers $a,b$ and $c$, set
$\mathbb{C}^{n}_{a,b,c}:=\left(\left(\mathbb{C}^{n}\right)^{\otimes a}\right)^{\oplus b}\otimes(\bigwedge^{n}\mathbb{C}^{n})^{\otimes-c}.$
Lemma 4.8.
Let $\rho:\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})\to\mathrm{GL}(W)$ be a representation that is homogeneous on the first factor, in the sense that the composition
[…]
is homogeneous. Then there exist $a,b$ and $c$ such that $W$ is a direct summand of $\mathbb{C}^{n}_{a,b,c}$ and $\rho$ extends to a representation
$\widetilde{\rho}:\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})\to\mathrm{GL}(\mathbb{C}^{n}_{a,b,c})$
such that the composition
$\widetilde{\rho}_{1}:\mathrm{GL}(n,\mathbb{C})\xrightarrow{\operatorname{id}\times 1}\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})\xrightarrow{\widetilde{\rho}}\mathrm{GL}(\mathbb{C}^{n}_{a,b,c})$
is induced by the standard representation of $\mathrm{GL}(n,\mathbb{C})$.
Proof.
This proof is modelled on [17, proof of 5.3, Proposition]. Let $M_{n}(\mathbb{C})$ be the ring of $(n\times n)$-matrices. Picking a basis of $W$, we may identify $\rho$ with a $(k\times k)$-matrix of regular functions in $\mathbb{C}[\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})]=\mathbb{C}[M_{n}(\mathbb{C})\times M_{n}(\mathbb{C})][\pi_{1}\circ\det^{-1},\pi_{2}\circ\det^{-1}]$, where $k=\dim W$ and $\pi_{i}:\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})\to\mathrm{GL}(n,\mathbb{C})$ is the projection onto the $i$-th factor. After replacing $\rho$ with $(\pi_{1}\circ\det)^{\otimes c}\otimes(\pi_{2}\circ\det)^{\otimes c}\otimes\rho$ for some $c$ big enough, we may assume that all these regular functions are in $\mathbb{C}[M_{n}(\mathbb{C})\times M_{n}(\mathbb{C})]$, so $\rho$ extends to a multiplicative map
$\rho:M_{n}(\mathbb{C})\times M_{n}(\mathbb{C})\to M_{k}(\mathbb{C}).$
The left $\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})$-action on $M_{n}(\mathbb{C})\times M_{n}(\mathbb{C})$ given by $g*A:=A\cdot g^{-1}$ provides $\mathbb{C}[M_{n}(\mathbb{C})\times M_{n}(\mathbb{C})]$ with the $\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})$-module structure
$(g*f)(A):=f(g^{-1}*A),$
where $g\in\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})$, $f\in\mathbb{C}[M_{n}(\mathbb{C})\times M_{n}(\mathbb{C})]$ and $A\in M_{n}(\mathbb{C})\times M_{n}(\mathbb{C})$. Given $\nu\in W^{*}$, define the linear map
$\varphi_{\nu}:W\to\mathbb{C}[M_{n}(\mathbb{C})\times M_{n}(\mathbb{C})];\,w\mapsto\left(A\mapsto\nu(\rho(A)(w))\right).$
It is straightforward to check that $\varphi_{\lambda}$ is a map of $\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})$-modules. Choosing a basis $\nu_{1},\dots,\nu_{k}$ of $W^{*}$, we obtain an injective map of $\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})$-modules
(20) $\varphi:W\hookrightarrow\mathbb{C}[M_{n}(\mathbb{C})\times M_{n}(\mathbb{C})]^{\oplus k};\,w\mapsto(\varphi_{\nu_{1}}(w),\dots,\varphi_{\nu_{k}}(w)).$
Now note that
$\mathbb{C}[M_{n}(\mathbb{C})\times M_{n}(\mathbb{C})]$ $=\mathbb{C}[M_{n}(\mathbb{C})]\otimes\mathbb{C}[M_{n}(\mathbb{C})]$
$=\operatorname{Sym}^{*}\left(M_{n}(\mathbb{C})^{*}\right)\otimes\operatorname{Sym}^{*}\left(M_{n}(\mathbb{C})^{*}\right)$
$\cong\operatorname{Sym}^{*}\left((\mathbb{C}^{n})^{\oplus n*}\right)\otimes\operatorname{Sym}^{*}\left((\mathbb{C}^{n})^{\oplus n*}\right)$
$\subset\bigoplus_{d_{1},d_{2}\geq 0}\left((\mathbb{C}^{n})^{\oplus n*}\right)^{\otimes d_{1}}\otimes\left((\mathbb{C}^{n})^{\oplus n*}\right)^{\otimes d_{2}}$
as $\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})$-modules. Combining this with (20), and using the fact that $W$ has finite dimension, we find an injective map of $\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})$-modules
$W\hookrightarrow\left(\bigoplus_{d\geq d_{1},d_{2}\geq 0}\left((\mathbb{C}^{n})^{n*}\right)^{\otimes d_{1}}\otimes\left((\mathbb{C}^{n})^{n*}\right)^{\otimes d_{2}}\right)^{\oplus k},$
for a big enough positive integer $d$. Since $\rho_{1}$ is homogeneous, $W$ is the degree $d_{1}$ part for some $d_{1}$, i.e.
$W\hookrightarrow\left(\bigoplus_{d\geq d_{2}\geq 0}\left((\mathbb{C}^{n})^{n*}\right)^{\otimes d_{1}}\otimes\left((\mathbb{C}^{n})^{n*}\right)^{\otimes d_{2}}\right)^{\oplus k}.$
Moreover, the composition
$\mathrm{GL}(n,\mathbb{C})\xrightarrow{\operatorname{id}\times 1}\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})\to\mathrm{GL}\left(\left(\bigoplus_{d\geq d_{2}\geq 0}\left((\mathbb{C}^{n})^{n*}\right)^{\otimes d_{1}}\otimes\left((\mathbb{C}^{n})^{n*}\right)^{\otimes d_{2}}\right)^{\oplus k}\right)$
is isomorphic to the representation
$\mathrm{GL}(n,\mathbb{C})\to\mathrm{GL}\left(\left((\mathbb{C}^{n})^{\otimes a}\right)^{\oplus b}\right)$
induced by the standard representation, for some positive integers $a$ and $b$. The proof is finished if we note that $\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})$-submodules and direct summands are the same, since $\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})$ is reductive. ∎
Consider the left linear $\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})$-action on $\mathbb{C}^{n}\otimes\mathbb{C}^{n*}$, given by
$\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})\times(\mathbb{C}^{n}\otimes\mathbb{C}^{n*})\to(\mathbb{C}^{n}\otimes\mathbb{C}^{n*});\,(g_{1},g_{2})\cdot(v,\nu)=(g_{2}\cdot v,g_{1}\cdot\nu),$
where the $\mathrm{GL}(n,\mathbb{C})$-actions on $\mathbb{C}^{n}$ and $\mathbb{C}^{n*}$ are the standard representation and the dual of the standard representation, respectively. This is homogeneous on both $\mathrm{GL}(n,\mathbb{C})$-factors. It induces a representation
$\rho:\mathrm{GL}(n,\mathbb{C})\times\mathrm{GL}(n,\mathbb{C})\to\mathrm{GL}\left(\mathcal{S}ym^{d}(\mathbb{C}^{n}\otimes\mathbb{C}^{n*})^{G}\right)$
which is also homogeneous on both factors, in the sense that the compositions
[…]
and
[…]
are both homogeneous. By Lemma 4.8, there exist integers $a,b$ and $c$ and a surjective homomorphism of $\mathrm{GL}(n,\mathbb{C})$-representations
$\varphi_{a,b,c}:\mathbb{C}^{n}_{a,b,c}:=\left(\left(\mathbb{C}^{n}\right)^{\otimes a}\right)^{\oplus b}\otimes(\bigwedge^{n}\mathbb{C}^{n})^{\otimes-c}\to\operatorname{Sym}^{d}(\mathbb{C}^{n}\otimes\mathbb{C}^{n*})^{G},$
where $\mathrm{GL}(n,\mathbb{C})$ acts on $\operatorname{Sym}^{d}(\mathbb{C}^{n}\otimes\mathbb{C}^{n*})^{G}$ via $\rho_{1}$, and such that $\rho_{2}$ extends to a linear $\mathrm{GL}(n,\mathbb{C})$-action on $\mathbb{C}^{n}_{a,b,c}$ commuting with the $\mathrm{GL}(n,\mathbb{C})$-action induced by the standard representation. Setting $W=\mathbb{C}^{n}_{a,b,c}$, equipped with the representation $\rho_{a,b,c}:\mathrm{GL}(n,\mathbb{C})\to\mathrm{GL}(\mathbb{C}^{n}_{a,b,c})$ induced by the standard one and the right action
$\mathbb{C}^{n}_{a,b,c}\times\mathrm{GL}(n,\mathbb{C})\to\mathbb{C}^{n}_{a,b,c};\,(v,g)\mapsto v\cdot g:=\rho_{2}(g)^{-1}v,$
we have the necessary ingredients to make sense of Definition 4.1.
The surjection $\varphi_{a,b,c}$ induces a surjective homomorphism of sheaves
(21) $\left(E^{\otimes a}\right)^{\oplus b}\otimes(\det E)^{\otimes-c}\to\mathcal{S}ym^{d}(\mathcal{O}_{X}^{\oplus n}\otimes E^{*})^{,}$
Therefore, the reduction of structure group $h$ is completely determined by the homomorphism
[…]
where $E(\mathbb{C}^{n}_{a,b,c})$ is the associated vector bundle via the representation $\rho_{a,b,c}:\mathrm{GL}(n,\mathbb{C})\to\mathrm{GL}(\mathbb{C}^{n}_{a,b,c})$ induced by the standard one. By (18), $\sigma$ is not the zero section. Moreover, since $(E,h,f)$ is a pseudo-equivariant $G$-bundle, (8) holds, which implies (11). Therefore, $(E,\sigma,f)$ is a pseudo-equivariant $\rho_{a,b,c}$-pair.
Summing up, the following holds.
Proposition 4.9.
There is an embedding of the category of pseudo-equivariant $G$-bundles into the category of pseudo-equivariant $\rho_{a,b,c}$-pairs sending $(E,h,f)$ to $(E,\sigma,f)$, where
(22) $\sigma:=h\circ\varphi_{a,b,c}:E(\mathbb{C}^{n}_{a,b,c}):=\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},E)\times_{\rho_{a,b,c}}\mathbb{C}^{n}_{a,b,c}\to\mathcal{O}_{X}.$
Remark 4.10.
Due to Remark 3.2, one may think that it is enough to consider $\rho_{a,b}$-pairs, rather than $\rho_{a,b,c}$-pairs, i.e. set $c=0$. However, the isomorphism $\det E\cong\mathcal{O}_{X}$ is not canonical, which is relevant when constructing the moduli space.
4.3. Semistable pseudo-equivariant $G$-bundles as semistable pseudo-equivariant pairs
This section is devoted to the proof of the following result, which compares the stability notions for pseudo-equivariant $G$-bundles and pseudo-equivariant $\rho_{a,b,c}$-pairs.
Proposition 4.11.
For $\chi\gg\delta\gg 0$, the following holds: given a pseudo-equivariant $G$-bundle $(E,h,f)$, the corresponding pseudo-equivariant pair $(E,\sigma,f)$ is $(\delta,\chi)$-(semi)stable if and only if $(E,h,f)$ is (semi)stable.
We break the proof in several lemmas.
Lemma 4.12.
Fix a pseudo-equivariant $G$-bundle $(E,h,f)$ with associated pseudo-equivariant pair $(E,\sigma,f)$. Then $\mu(E_{\bullet},\alpha_{\bullet},f)\geq 0$ for any weighted filtration, with $\mu(E_{\bullet},\alpha_{\bullet},f)=0$ if and only if the filtration $E_{\bullet}$ is $f$-invariant.
Proof.
Let $\lambda$ be a 1-parameter subgroup of $\mathrm{GL}(n,\mathbb{C}(X))$ determining $(E_{\bullet},\alpha_{\bullet})$, where $\mathbb{C}(X)$ is the field of functions on $X$. It is clear that the weights appearing in the decomposition of $f|_{\eta}$ under the action of $\lambda$ are precisely the negative of the weights of $f^{*-1}|_{\eta}$. By Remark 4.3 and Definition (12), $\mu(E_{\bullet},\alpha_{\bullet},f)\geq 0$. Moreover, $\mu(E_{\bullet},\alpha_{\bullet},f)=0$ if and only if all the weights of $f|_{\eta}$ are 00, which happens if and only if $E_{\bullet}$ is $f$-invariant. ∎
Following [21], we introduce a concept closely related to $(\delta,\chi)$-(semi)stability for pseudo-equivariant pairs.
Definition 4.13.
A pseudo-equivariant pair $(E,\sigma,f)$ is called asymptotically (semi)stable if both $\mu(E_{\bullet},\alpha_{\bullet},\sigma)$ and $\mu(E_{\bullet},\alpha_{\bullet},f)$ are non-negative for every weighted filtration $(E_{\bullet},\alpha_{\bullet})$ and, moreover, every proper weighted filtration such that $\mu(E_{\bullet},\alpha_{\bullet},\sigma)=\mu(E_{\bullet},\alpha_{\bullet},f)=0$ satisfies $M(E_{\bullet},\alpha_{\bullet})\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$.
Lemma 4.14.
A pseudo-equivariant $G$-bundle $(E,h,f)$ is (semi)stable if and only if the corresponding pseudo-equivariant pair $(E,\sigma,f)$ is asymptotically (semi)stable.
Proof.
First assume that $(E,\sigma,f)$ is asymptotically (semi)stable. Fix a $\Gamma$-invariant maximal compact subgroup of $G$ with Lie algebra $\mathfrak{k}\subset\mathfrak{gl}(n,\mathbb{C})$. For every $s\in i\mathfrak{k}^{\Gamma}$, we may define parabolic subgroups $\widetilde{P}_{s}\subset\mathrm{GL}(n,\mathbb{C})$ and $P_{s}:=\widetilde{P}_{s}\cap G$, where
(23) $\widetilde{P}_{s}:=\{g\in\mathrm{GL}(n,\mathbb{C})\;:\;e^{ts}ge^{-ts}\,\text{remains bounded as}\;t\to\infty\}.$
Let $Q:=\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},E)$ be the bundle of frames of $E$ — a $\mathrm{GL}(n,\mathbb{C})$-bundle —, and let $Q_{G}$ be the $G$-bundle associated to $h$. Let $\tau\in H^{0}(X,Q_{G}/P_{s})^{\Gamma}$ be a $\Gamma$-invariant reduction of structure group with corresponding $P_{s}$-principal bundle $Q_{\tau}$. We have to prove that $\deg Q_{G}(\tau,s)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$.
The extension of structure group of $Q_{\tau}$ to $\widetilde{P}_{s}$ provides a reduction of structure group of the bundle of frames of $E$ to $\widetilde{P}_{s}$, which we denote by $\widetilde{\tau}\in H^{0}(X,Q/\widetilde{P}_{s})$, with corresponding $\widetilde{P}_{s}$-bundle $Q_{\widetilde{\tau}}$. This determines a filtration $E_{\bullet}$ of $E$, and the 1-parameter subgroup
$\mathbb{C}^{*}\to\mathrm{GL}(n,\mathbb{C});\,t\mapsto e^{ts}$
determines a corresponding system of weights $\alpha_{\bullet}$, thus defining a weighted filtration $(E_{\bullet},\alpha_{\bullet})$ such that $M(E_{\bullet},\alpha_{\bullet})=\deg Q_{G}(\tau,s)$. By assumption, $M(E_{\bullet},\alpha_{\bullet})\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$ as long as $\mu(E_{\bullet},\alpha_{\bullet},\sigma)=\mu(E_{\bullet},\alpha_{\bullet},f)=0$.
Let $\eta$ be the generic point of $X$. Then there is an isomorphism $Q_{G}|_{\eta}\cong G\times_{\mathbb{C}}\mathbb{C}(X)$ sending $Q_{\tau}|_{\eta}$ to $P_{s}\times_{\mathbb{C}}\mathbb{C}(X)$, where $\mathbb{C}(X)$ is the field of functions of $X$. Under this isomorphism $Q_{\tau}|_{\eta}$ is determined by the one parameter subgroup generated by the exponential of $s\otimes 1\in\mathfrak{p}_{s}\otimes\mathbb{C}(X)$. This extends to an isomorphism $Q|_{\eta}\cong\mathrm{GL}(n,\mathbb{C}(X))$ such that $Q_{\widetilde{\tau}}|_{\eta}$ is also determined by $s\otimes 1$. Hence, we may choose the 1-parameter subgroup $\lambda$ of $\mathrm{GL}(E|_{\eta})$ determining $(E_{\bullet},\alpha_{\bullet})$ to be generated by the exponential of $s\otimes 1$. The $\mathbb{C}(X)$-action defined by $\lambda$ preserves $Q_{G}|_{\eta}$ — in other words, it preserves $h|_{\eta}$ and $\sigma|_{\eta}$, hence $\mu(\sigma,\lambda)=0$.
Finally, the fact that $\tau$ is $\Gamma$-invariant implies that $\widetilde{\tau}$ is $\Gamma$-invariant and so is $\lambda$, hence we conclude that $\mu(f|_{\eta},\lambda)=0$ by Lemma 4.12. Therefore, $\mu(E_{\bullet},\alpha_{\bullet},\sigma)=\mu(E_{\bullet},\alpha_{\bullet},f)=0$ and so $\deg Q_{G}(\tau,s)=M(E_{\bullet},\alpha_{\bullet})\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$ as required.
To prove the other direction, let $(E,h,f)$ be a (semi)stable pseudo-equivariant $G$-bundle. Let $(E_{\bullet},\alpha_{\bullet})$ be a proper weighted filtration of $E$, assume without loss that $\alpha_{\bullet}$ is integral and let $\lambda:\mathbb{G}_{m}(\mathbb{C}(X))\to\mathrm{GL}(n,\mathbb{C}(X))$ be a 1-parameter subgroup determining $(E_{\bullet},\alpha_{\bullet})$. By Proposition 2.5, the underlying $G$-bundle $(E,h)$ is semistable. By [21, Corollary 2.4.4.6], for $\delta\gg 0$ the underlying $\rho_{a,b,c}$-swamp $(E,\mathcal{O}_{X},\sigma)$ is $\delta$-semistable — in the sense of [21, Section 2.3.2], see Remark 4.5 —, hence $\mu(E_{\bullet},\alpha_{\bullet},\sigma)\geq 0$ for $\delta\gg 0$ by [21, Proposition 2.3.6.5]. Moreover $\mu(E_{\bullet},\alpha_{\bullet},f)\geq 0$ by Lemma 4.12, so the first condition in Definition 4.13 holds.
Now assume that $\mu(E_{\bullet},\alpha_{\bullet},\sigma)=\mu(E_{\bullet},\alpha_{\bullet},f)=0$, i.e. $\mu(\sigma|_{\eta},\lambda)=\mu(f|_{\eta},\lambda)=0$. Then $h|_{\eta}$ is invariant under the $\mathbb{C}^{*}$-action given by $\lambda$. Therefore, $\lambda$ determines a parabolic subgroup $P$ of $G$ and a reduction of structure group $\tau\in H^{0}(X,Q_{G}/P)$ such that the reduction of structure group of $Q=\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},E)$ determined by $E_{\bullet}$ is an extension of structure group of $Q_{\tau}$. By [21, Proposition 2.4.9.1] we may also choose $s\in\mathfrak{g}$ such that $P=P_{s}$ and the 1-parameter subgroup of $\mathrm{GL}(n,\mathbb{C}(X))$ generated by the exponential of $s\otimes 1$ determines the weighted filtration $(E_{\bullet},r\alpha_{\bullet})$ for some $r\in\mathbb{Q}_{\geq 0}$, so that $\deg Q_{G}(\tau,s)=rM(E_{\bullet},\alpha_{\bullet})$. Note that $s\neq 0$, since $E_{\bullet}$ is a proper filtration of $E$. Therefore, since $G$ is semisimple, $P_{s}\subset G$ is a proper parabolic subgroup.
It is left to show that $\deg Q_{G}(\tau,s)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$. Recall from Section 2.1 that a $\Gamma$-invariant compact subgroup $K$ has been previously fixed in order to define the stability notions for twisted equivariant $G$-bundles — or pseudo-equivariant $G$-bundles. We may also choose a $\Gamma$-invariant maximal torus $T\subset G$ such that $T\cap K$ is a maximal torus of $K$. First note that, given $g\in G$, we may use the $G$-action on $Q$ to define a reduction of structure group $\tau g\in H^{0}(X,Q/g^{-1}P_{s}g)$ to $g^{-1}P_{s}g$, with corresponding weighted filtration $(E_{\bullet}g,\alpha_{\bullet})$. The induced pseudo-equivariant action $g^{-1}fg$ satisfies $\mu(E_{\bullet}g,\alpha_{\bullet},g^{-1}fg)=\mu(E_{\bullet},\alpha_{\bullet},f)=0$. Therefore, we may assume after conjugating by an element $g\in G$ that $s\in i\mathfrak{k}$ and the corresponding antidominant character $\chi_{s}$ is in a $\Gamma$-invariant Weyl chamber.
Let $\widetilde{\tau}\in H^{0}(X,Q/\widetilde{P}_{s})$ be the reduction of structure group given by extending the structure group of $Q_{\tau}$ to $\widetilde{P}_{s}$. By the assumption $\mu(E_{\bullet},\alpha_{\bullet},f)=0$ and Lemma 4.12, we know that $Q_{\tau^{\prime}}$ is invariant by the equivariant action given by $f$, i.e. $f_{\gamma}$ sends $Q_{\tau^{\prime}}$ to $\gamma^{*}Q_{\tau^{\prime}}\subset\gamma^{*}Q$. By (8) $f_{\gamma}(Q_{\tau})\subset\gamma^{*}Q_{\tau}(1,\gamma)^{-1}$, which implies that the intersection of $\widetilde{P}_{s}$ with $G(1,\gamma)^{-1}$ inside of $\mathrm{GL}(n,\mathbb{C})$ is non-empty for each $\gamma\in\Gamma$. Therefore, the subgroup $\overline{P}_{s}:=\widetilde{P}_{s}\cap(G\times_{(\theta,c)}\Gamma)\subset G\times_{(\theta,c)}\Gamma\subset\mathrm{GL}(n,\mathbb{C})$, which has connected component of the identity $P_{s}$, intersects all the connected components of $G\times_{(\theta,c)}\Gamma$. Since $P_{s}$ is a normal subgroup of $\overline{P}_{s}$, there exists $g_{\gamma}\in G(1,\gamma)$ — namely, any element $g_{\gamma}\in\overline{P}_{s}\cap G(1,\gamma)$ — such that $g_{\gamma}P_{s}g_{\gamma}^{-1}=P_{s}$, for each $\gamma\in\Gamma$. Let $L_{s}$ be the levi subgroup of $P_{s}$ defined by (6), and let
(24) $R_{s}:=\{g\in G\;:\;\lim_{t\to\infty}e^{ts}ge^{-ts}=1\},$
be the radical of $P_{s}$ — that is, its maximal solvable normal subgroup. Since every two Levi subgroups of $P_{s}$ are conjugate we may assume, after conjugating $g_{\gamma}$ by an element of $P_{s}$, that $\operatorname{Int}_{g_{\gamma}}$ preserves $L_{s}$. Since every two maximal tori in $L_{s}$ are conjugate to each other, after conjugating $g_{\gamma}$ further by an element of $L_{s}$ we may assume that $\operatorname{Int}_{g_{\gamma}}$ preserves the maximal torus $T\cap L_{s}$. Note that $\operatorname{Int}_{g_{\gamma}}$ preserves $R_{s}$, and in particular its maximal torus $T\cap R_{s}$. As $T\cap R_{s}$ and $T\cap L_{s}$ generate $T$, we may assume that $\operatorname{Int}_{g_{\gamma}}$ preserves $T$. Since $\operatorname{Int}_{g_{\gamma}}$ and $\operatorname{Int}_{(1,\gamma)}=\theta_{\gamma}$ are in the same outer class, we conclude that $\operatorname{Int}_{g_{\gamma}}=\theta_{\gamma}\operatorname{Int}_{g_{0,\gamma}}$ for some $g_{0,\gamma}\in G$ preserving $T$.
Therefore, $\theta_{\gamma}\operatorname{Int}_{g_{0,\gamma}}$ preserves $P_{s}$. As $\theta_{\gamma}$ preserves the Weyl chamber where $\chi_{s}$ lies, the element of the Weyl group defined by $\operatorname{Int}_{g_{0,\gamma}}$ is trivial. In particular, the $\theta_{\gamma}\operatorname{Int}_{g_{0,\gamma}}$-invariance of $T\cap L_{s}$ and $T\cap R_{s}$ implies that both $T\cap L_{s}$ and $T\cap R_{s}$ are $\theta_{\gamma}$-invariant, which in turn implies that $L_{s}$ and $P_{s}$ are $\theta_{\gamma}$-invariant for each $\gamma\in\Gamma$. Therefore $s_{\Gamma}:=\sum_{\gamma\in\Gamma}\theta_{\gamma}(s)$, which is non-zero because sums of non-zero vectors in a Weyl chamber are never zero, is in $i\mathfrak{k}^{\Gamma}$ and satisfies $P_{s_{\Gamma}}=P_{s}$. By (semi)stability of $(E,h,f)$ we know that $\deg Q_{G}(\tau,s_{\Gamma})\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$. Moreover, (8) implies $F_{A}=\gamma^{*}\theta_{\gamma}^{-1}(F_{A})$ for each $\gamma\in\Gamma$, where $F_{A}$ is the curvature of a connection $A$ on $K\cap L_{s}$. Therefore,
$\deg Q_{G}(\tau,s_{\Gamma})$ $=\frac{i}{2\pi}\int_{X}\langle s_{\Gamma},F_{A}\rangle$
$=\frac{i}{2\pi}\sum_{\gamma\in\Gamma}\int_{X}\langle\theta_{\gamma}(s),F_{A}\rangle$
$=\frac{i}{2\pi}\sum_{\gamma\in\Gamma}\int_{X}\langle\theta_{\gamma}(s),\gamma^{*-1}\theta_{\gamma}F_{A}\rangle$
$=\frac{i}{2\pi}\sum_{\gamma\in\Gamma}\int_{X}\langle\theta_{\gamma}(s),\theta_{\gamma}F_{A}\rangle$
$=\frac{i}{2\pi}\sum_{\gamma\in\Gamma}\int_{X}\langle s,F_{A}\rangle$
$=|\Gamma|\deg Q_{G}(\tau,s),$
[…]
Lemma 4.15.
For $\chi\gg\delta\gg 0$, a pseudo-equivariant pair $(E,\sigma,f)$ coming from a pseudo-equivariant $G$-bundle $(E,h,f)$ is asymptotically (semi)stable only if it is $(\delta,\chi)$-(semi)stable.
Proof.
Assume that $(E,\sigma,f)$ is asymptotically (semi)stable. By Lemma 4.14, the corresponding pseudo-equivariant $G$-bundle is (semi)stable, and so is the underlying $G$-bundle by Proposition 2.5. By [21, Corollary 2.4.4.6], for $\delta\gg 0$ the corresponding $\rho_{a,b,c}$-swamp $(E,\mathcal{O}_{X},\sigma)$ is $\delta$-semistable. By [21, Theorem 2.3.4.1] and Remark 3.2, there exists a constant $C$, independent of $(E,\sigma,f)$, such that
(25) $\deg(F)/\operatorname{rk}(F)\leq C.$
By Lemma 4.6, $(\delta,\chi)$-(semi)stability only has to be verified on weighted filtrations $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$, for some finite set $S$. We may assume that $S$ only features integral weights. Using (25), we may find a negative number $-\epsilon_{0}$ such that $M(E_{\bullet},\alpha_{\bullet})>-\epsilon_{0}$ for such weighted filtrations. Choose $\delta\geq\epsilon_{0}$. Using [21, Lemma 1.5.1.41], we may find a negative number $-\epsilon_{1}$ such that $M(E_{\bullet},\alpha_{\bullet})+\delta\mu(E_{\bullet},\alpha_{\bullet},\sigma)>-\epsilon_{1}$, for every weighted filtration $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$. Choose $\chi\geq\epsilon_{1}$. Then, for such weighted filtrations, since the weights are integral we find
$M(E_{\bullet},\alpha_{\bullet})+\delta\mu(E_{\bullet},\alpha_{\bullet},\sigma)+\chi\mu(E_{\bullet},\alpha_{\bullet},f)>0$
whenever $\mu(E_{\bullet},\alpha_{\bullet},f)$ or $\mu(E_{\bullet},\alpha_{\bullet},\sigma)>0$. If, on the other hand, $\mu(E_{\bullet},\alpha_{\bullet},f)=\mu(E_{\bullet},\alpha_{\bullet},\sigma)=0$ — this is the only remaining case, since $\mu(E_{\bullet},\alpha_{\bullet},f),\mu(E_{\bullet},\alpha_{\bullet},\sigma)\geq 0$ by asymptotic (semi)stability —, then
$M(E_{\bullet},\alpha_{\bullet})+\delta\mu(E_{\bullet},\alpha_{\bullet},\sigma)+\chi\mu(E_{\bullet},\alpha_{\bullet},f)=M(E_{\bullet},\alpha_{\bullet})\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$
by Definition 4.13, as required. ∎
Lemma 4.16.
A pseudo-equivariant $G$-bundle $(E,h,f)$ is (semi)stable if the corresponding pseudo-equivariant pair $(E,\sigma,f)$ is $(\delta,\chi)$-(semi)stable.
Proof.
Note that, given a weighted filtration $(E_{\bullet},\alpha_{\bullet})$ such that
(26) $\mu(E_{\bullet},\alpha_{\bullet},\sigma)=\mu(E_{\bullet},\alpha_{\bullet},f)=0,$
then
$M(E_{\bullet},\alpha_{\bullet})+\delta\mu(E_{\bullet},\alpha_{\bullet},\sigma)+\chi\mu(E_{\bullet},\alpha_{\bullet},f)=M(E_{\bullet},\alpha_{\bullet})$
for every $\delta$ and $\chi$. Therefore, if $(E,\sigma,f)$ is $(\delta,\chi)$-(semi)stable, then (26) implies that $M(E_{\bullet},\alpha_{\bullet})\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$. Using this fact, the proof is identical to the first part of the proof of Lemma 4.14. ∎
Proposition 4.11 follows from Lemmas 4.14, 4.15 and 4.16.
5. The parameter space of twisted equivariant $G$-bundles
Fix a compact Riemann surface $X$, a finite group $\Gamma$ of complex automorphisms of $X$, a connected semisimple complex Lie group $G$ with center $Z$, a homomorphism $\theta:\Gamma\to\operatorname{Aut}(G)$ and a 2-cocycle $c\in Z^{2}_{\theta}(\Gamma,Z)$. By Propositions 3.3, 4.9 and 4.11, if we construct the moduli space of $(\delta,\chi)$-semistable pseudo-equivariant $\rho_{a,b,c}$-pairs then the moduli space of semistable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles will be embedded in it for $\chi\gg\delta\gg 0$.
Throughout this section, given two spaces $Y$ and $Z$, we denote by $\pi_{Y}:Y\times Z\to Y$ the projection.
5.1. Bounded families of vector bundles and the Quot Scheme
Recall that a family of vector bundles $\mathfrak{G}$ is bounded if there exists a scheme $S$ of finite type over $\mathbb{C}$ and a vector bundle $E_{S}$ over $S\times_{\mathbb{C}}X$ such that, for every vector bundle $E$ in $\mathfrak{G}$, there exists $s\in S$ such that $E_{S}|_{\{s\}\times X}\cong E$.
Assume that $\chi\gg\delta\gg 0$, so that we may use Proposition 4.11. By [4, Proposition 4.1] the family of semistable twisted equivariant $G$-bundles over $X$ has an underlying family of semistable $G$-bundles, whose associated family $\mathfrak{G}$ of vector bundles of rank $n$ and trivial determinant — obtained via extension of structure group to $\mathrm{GL}(n,\mathbb{C})$ — is bounded by [21, Corollary 2.4.4.6, Proposition 2.2.3.7 and Theorem 2.3.4.1]. Let $\mathcal{O}_{X}(1)$ be a $\Gamma$-invariant ample line bundle on $X$; this can be achieved by taking any ample line bundle and tensoring it by its pullbacks with respect to the elements of $\Gamma$. Set $E(m):=E\otimes\mathcal{O}_{X}(1)^{\otimes m}$, where $E$ is a vector space or a vector bundle over $X$. By [21, Corollary 2.2.3.4] and Remark 3.2, we know that there exist a complex vector space $V$ and a natural number $m$ satisfying the following.
Proposition 5.1.
Every vector bundle $E$ of rank $n$ and trivial determinant associated to a semistable twisted equivariant $G$-bundle is a quotient $V(-m)\to E$ over $X$ such that the induced homomorphism $V\to H^{0}(E(m))$ on spaces of global sections is an isomorphism.
There is a Grothendieck’s Quot Scheme $\mathfrak{Q}$ equipped with a family of vector bundles $\mathfrak{F}$ over $\mathfrak{Q}\times X$ of rank $n$ and trivial determinant over $X$, and a quotient homomorphism $q:\pi_{X}^{*}V(-m)\to\mathfrak{F}$ of sheaves over $\mathfrak{Q}\times X$. The quasiprojective variety $\mathfrak{Q}$ is a fine moduli space for equivalence classes of vector bundles of rank $n$ and trivial determinant which are quotients of $V(-m)$ over $X$, where the equivalence relation is given by having the same kernel. By Proposition 5.1, for every $(\delta,\chi)$-semistable $\rho_{a,b,c}$-pair $(E,\sigma,f)$ coming from a twisted equivariant $G$-bundle, there exists $t\in\mathfrak{Q}$ such that $\mathfrak{F}|_{\{t\}\times X}\cong E$.
5.2. The parameter space of $G$-bundles
In [21] the moduli space of holomorphic principal $G$-bundles is constructed using $\rho$-swamps, which are triples $(E,L,\varphi)$ made of a vector bundle $E$, a line bundle $L$ and a surjection $\varphi:E(W)\to L$, where $\rho:\mathrm{GL}(n,\mathbb{C})\to W$ is some representation and $E(W)=\mathcal{H}om(\mathcal{O}_{X}^{\oplus n},E)\times_{\rho}W$ is the associated vector bundle. Indeed, if we ignore the pseudo-equivariant structure in Section 4.2, we obtain an embedding of the category of $G$-bundles into the category of $\rho_{a,b,c}$-pairs $(E,\sigma)$ consisting of a vector bundle $E$ of rank $n$ and degree 00, and a homomorphism $\sigma:E(\mathbb{C}^{n}_{a,b,c})\to\mathcal{O}_{X}$ — these are just $\Gamma$-pseudo-equivariant $\rho_{a,b,c}$-pairs for $\Gamma=1$. In what follows we build a parameter space for this pairs, following [21, Section 2.3].
Given a vector bundle $E$, we write $E_{a,b,c}:=(E^{\otimes a})^{\oplus b}\otimes(\det E)^{\otimes-c}$. Let $k$ be big enough so that $V_{k}:=\pi_{\mathfrak{Q}*}(\pi_{X}^{*}V(-m)_{a,b,c}\otimes\pi_{X}^{*}\mathcal{O}_{X}(ak))=\pi_{\mathfrak{Q}*}(\pi_{X}^{*}V_{a,b,c}\otimes\pi_{X}^{*}\mathcal{O}_{X}(a(k-m)))$ and $\pi_{\mathfrak{Q}*}\pi_{X}^{*}\mathcal{O}_{X}(ak)$ are vector bundles, where $\pi_{\mathfrak{Q}}:\mathfrak{Q}\times X\to\mathfrak{Q}$ is the projection. Let $\mathfrak{B}:=\mathbb{P}\left(\mathcal{H}om(V_{k},\pi_{\mathfrak{Q}*}\pi_{X}^{*}\mathcal{O}_{X}(ak))\right)$, which comes equipped with a projective bundle projection $p_{\mathfrak{B}}:\mathfrak{B}\to\mathfrak{Q}$. There is a tautological homomorphism
$(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\pi_{X}^{*}V(-m)_{a,b,c}\otimes(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\pi_{X}^{*}\mathcal{O}_{X}(ak)\to(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\pi_{X}^{*}\mathcal{O}_{X}(ak)\otimes\mathcal{O}_{\mathfrak{B}}(1)$
or, in other words, a homomorphism
$\varphi^{\prime}:(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\pi_{X}^{*}V(-m)_{a,b,c}\to\pi_{\mathfrak{B}}^{*}\mathcal{O}_{\mathfrak{B}}(1).$
Let $\mathfrak{F}_{a,b,c}:=(\mathfrak{F}^{\otimes a})^{\oplus b}\otimes(\det\mathfrak{F})^{\otimes-c}$ and consider the map
$q_{a,b,c}:(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\pi_{X}^{*}V(-m)_{a,b,c}\to(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\mathfrak{F}_{a,b,c}$
induced by $q$.
Remark 5.2.
Note that every $\rho_{a,b,c}$-pair providing a semistable $G$-bundle is isomorphic to
$((p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\mathfrak{F}_{a,b,c}|_{\{t\}\times X},\varphi|_{\{t\}\times X})$
for some $t\in\mathfrak{B}$. Indeed, for every vector bundle $E$ of rank $n$ and trivial determinant over $X$, given a quotient $V(-m)\to E$ inducing an isomorphism on global sections $V\xrightarrow{\sim}H^{0}(E(m))$, any homomorphism $\sigma:E_{a,b,c}\to\mathcal{O}_{X}$ lifts to a homomorphism $V(-m)_{a,b,c}\to\mathcal{O}_{X}$ via the composition
[…]
which induces a homomorphism
$V(-m)_{a,b,c}\cong V_{a,b,c}(-ma)\to\mathcal{O}_{X}.$
We may thus apply Proposition 5.1.
By [21, Proposition 2.3.5.1], the set of points $t\in\mathfrak{B}$ such that the restriction of $\varphi^{\prime}|_{\{t\}\times X}$ to the kernel of $q_{a,b,c}$ is zero is a closed subscheme ${\mathfrak{I}}_{1}\subset\mathfrak{B}$. On this subscheme, $\varphi^{\prime}$ induces a homomorphism
$\varphi:(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\mathfrak{F}_{a,b,c}|_{{\mathfrak{I}}_{1}}\to(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\pi_{X}^{*}\mathcal{O}_{X}|_{{\mathfrak{I}}_{1}}.$
Again by Proposition 2.3.5.1, we find a closed subscheme ${\mathfrak{I}}_{0}$ of ${\mathfrak{I}}_{1}$ where the restriction of $\varphi$ to the kernel of
$(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\mathfrak{F}_{a,b,c}\to\mathcal{S}ym^{*}(\mathcal{O}_{X}^{\oplus n}\otimes(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\mathfrak{F}^{*})^{G}$
is zero.
Finally, there is an open subvariety ${\mathfrak{I}}$ of ${\mathfrak{I}}_{0}$ consisting of the points $t\in{\mathfrak{I}}_{0}$ such that $\varphi|_{\{t\}\times X}$ induces a $G$-bundle, namely those inducing a section of $\mathcal{H}om(\mathcal{O}_{X}^{\oplus n},(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\mathfrak{F})\sslash G$ which actually lies in $\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\mathfrak{F})/G$.
We rename $\varphi$ to be its restriction to ×$\mathfrak{I}\times X$. Set $V_{\mathfrak{I}}:=V\otimes\mathcal{O}_{\mathfrak{I}\times X}$, $\mathfrak{F}_{\mathfrak{I}}:=(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\mathfrak{F}|_{{\mathfrak{I}}\times X}$ and $q_{\mathfrak{I}}:=(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}q|_{{\mathfrak{I}}\times X}$. We have obtained a universal family $(\mathfrak{F}_{\mathfrak{I}},\varphi)$ over ×${\mathfrak{I}}\times X$ parametrizing $G$-bundles over $X$ whose extension of structure group to $\mathrm{GL}(n,\mathbb{C})$ has associated vector bundle isomorphic to a quotient of $V(-m)$. Note that, for each $\gamma\in\Gamma$, there are alternative universal families $((\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}\mathfrak{F}_{\mathfrak{I}},(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}\varphi)$ and a quotient
$(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}q_{\mathfrak{I}}:(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}V_{\mathfrak{I}}(-m)\to(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}\mathfrak{F}_{\mathfrak{I}}.$
Remark 5.3.
Note that, given two $G$-bundles $(E,h)$ and $(E,h^{\prime})$ represented by the same point $t\in\mathfrak{B}$, the corresponding homomorphisms
$\left(V^{\otimes a}\right)^{\oplus b}\otimes(\det V)^{\otimes-c}\to\mathcal{O}_{X}$
— given by Section 4.2 and $q$ — are related by multiplication by an element of $\mathbb{C}^{*}$. In particular, they are in the same $\mathrm{GL}(V)$-orbit. We will see in Proposition 5.5 that this implies that $(E,h)\cong(E,h^{\prime})$. In other words, working with $\mathfrak{B}=\mathbb{P}\left(\mathcal{H}om(V_{k},\pi_{\mathfrak{Q}*}\pi_{X}^{*}\mathcal{O}_{X}(ak))\right)$ instead of $\mathcal{H}om(V_{k},\pi_{\mathfrak{Q}*}\pi_{X}^{*}\mathcal{O}_{X}(ak))$ does not lose information.
5.3. The parameter space of twisted equivariant pairs
With notation as in Section 5.2, let
$V_{\Gamma,\mathfrak{I}}:=\bigoplus_{\gamma\in\Gamma}\mathcal{H}om(V_{\mathfrak{I}},(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}V_{\mathfrak{I}})$
and
$V_{\Gamma,\mathfrak{I}}^{\vee}:=\bigoplus_{\gamma\in\Gamma}\mathcal{H}om(V^{*}_{\mathfrak{I}},(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}V^{*}_{\mathfrak{I}})$
Set $\mathfrak{C}:=\mathbb{P}(V_{\Gamma,\mathfrak{I}}\oplus V_{\Gamma,\mathfrak{I}}^{\vee})$, and let $p_{\mathfrak{C}}:\mathfrak{C}\to\mathfrak{I}$ be the bundle projection. We endow $\mathfrak{C}$ with the $\mathrm{GL}(V)$-left action induced by
$\mathrm{GL}(V)\times(V_{\Gamma,\mathfrak{I}}\oplus V_{\Gamma,\mathfrak{I}}^{\vee})\to V_{\Gamma,\mathfrak{I}}\oplus V_{\Gamma,\mathfrak{I}}^{\vee};\,g\cdot(\bigoplus_{\gamma\in\Gamma}\psi^{\prime}_{\gamma},\bigoplus_{\gamma\in\Gamma}\psi^{\prime\prime}_{\gamma}):=(\bigoplus_{\gamma\in\Gamma}g\psi^{\prime}_{\gamma}g^{-1},\bigoplus_{\gamma\in\Gamma}g^{t-1}\psi^{\prime\prime}_{\gamma}g^{t}).$
This comes equipped with a tautological section $\psi^{\prime}\oplus\psi^{\prime\prime}\in H^{0}((p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}(V_{\Gamma,\mathfrak{I}}\oplus V_{\Gamma,\mathfrak{I}}^{\vee})\otimes\mathcal{O}_{\mathfrak{C}}(1))$. Let $\psi^{\prime}\in H^{0}((p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}V_{\Gamma,\mathfrak{I}}\otimes\mathcal{O}_{\mathfrak{C}}(1))$ be its projection onto the first summand, and write
$\psi^{\prime}=\bigoplus_{\gamma\in\Gamma}\psi^{\prime}_{\gamma}:(p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}V_{\mathfrak{I}}(-m)\to\bigoplus_{\gamma\in\Gamma}(p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}V_{\mathfrak{I}}(-m)\otimes\mathcal{O}_{\mathfrak{C}}(1).$
Note that the operations of pulling back by $\operatorname{id}_{\mathfrak{I}}\times\gamma$ and tensoring by $\mathcal{O}_{X}(-m)$ commute by $\Gamma$-invariance of $\mathcal{O}_{X}(1)$, so we don’t need to especify the order.
Let $\mathfrak{C}_{1}\subset\mathfrak{C}$ be the subset of points $t$ such that the restriction of $\psi^{\prime}_{\gamma}$ to $\{t\}\times X$ is an isomorphism for every $\gamma\in\Gamma$. Since the degrees of a line bundle over $X$ and its pullback by an element of $\Gamma$ are equal, this is the complement of the locus of points $t\in\mathfrak{C}$ such that the induced homomorphisms
$\det\psi^{\prime}_{\gamma}:\det(p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}V_{\mathfrak{I}}(-m)\to\det(p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}V_{\mathfrak{I}}(-m)\otimes\mathcal{O}_{\mathfrak{C}}(1)$
are zero on $\{t\}\times X$, for each $\gamma\in\Gamma$. By [21, Proposition 2.3.5.1] this locus is closed, so $\mathfrak{C}_{1}$ is open in $\mathfrak{C}$.
Let $\psi_{0}$ be the image of $\psi^{\prime}$ by the homomorphism
$(p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}V_{\Gamma,\mathfrak{I}}\to(p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}\bigoplus_{\gamma\in\Gamma}\mathcal{H}om(\ker q_{\mathfrak{I}},(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}\mathfrak{F}_{\mathfrak{I}})$
of vector bundles over ×${\mathfrak{I}}\times X$ given by (the pullback of) restricting each homomorphism $V_{\mathfrak{I}}(-m)\to(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}V_{\mathfrak{I}}(-m)$ to $\ker q_{\mathfrak{I}}$ and composing with $(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}q_{\mathfrak{I}}$. The locus of $\mathfrak{C}$ where $\psi_{0}$ is zero is a closed subscheme $\mathfrak{C}_{2}$ of $\mathfrak{C}$ by [21, Proposition 2.3.5.1]. This is precisely the subvariety where $\psi^{\prime}$ descends to a section $\psi$ of
$\mathfrak{F}_{\Gamma,\mathfrak{C}}:=(p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}\bigoplus_{\gamma\in\Gamma}\mathcal{H}om(\mathfrak{F}_{\mathfrak{I}},(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}\mathfrak{F}_{\mathfrak{I}}).$
Again by [21, Proposition 2.3.5.1], the locus of $\mathfrak{C}$ where the image of $\psi^{\prime}$ by the homomorphism
$p_{\mathfrak{C}}^{*}V_{\Gamma,\mathfrak{I}}\to p_{\mathfrak{C}}^{*}\bigoplus_{(\gamma,\gamma^{\prime})\in\Gamma\times\Gamma}\mathcal{H}om(V_{\mathfrak{I}},(\operatorname{id}_{\mathfrak{I}}\times\gamma\gamma^{\prime})^{*}V_{\mathfrak{I}});$
$p_{\mathfrak{C}}^{*}(f_{\gamma})_{\gamma\in\Gamma}\mapsto\bigoplus_{\gamma\in\Gamma}p_{\mathfrak{C}}^{*}((\gamma^{*}f_{\gamma^{\prime}})f_{\gamma}-f_{\gamma\gamma^{\prime}})$
is 00 is a closed subscheme $\mathfrak{C}_{3}\subset\mathfrak{C}$. We also consider the closed subscheme $\mathfrak{C}_{4}$ where $\psi^{\prime}_{1}=\operatorname{id}$.
Set $\mathfrak{F}_{a,b,c,\mathfrak{J}}:=(p_{\mathfrak{B}}\times\operatorname{id}_{X})^{*}\mathfrak{F}_{a,b,c}|_{\mathfrak{J}}$. Let $\mathfrak{C}_{5}\subset\mathfrak{C}_{2}$ be the closed subscheme where the image of $p_{\mathfrak{C}}^{*}\varphi$ by the homomorphism
$\mathcal{H}om((p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}\mathfrak{F}_{a,b,c,\mathfrak{J}},\mathcal{O}_{\mathfrak{C}\times X})\to\bigoplus_{\gamma\in\Gamma}\mathcal{H}om((p_{\mathfrak{C}}\times\gamma)^{*}\mathfrak{F}_{a,b,c,\mathfrak{J}},\mathcal{O}_{\mathfrak{C}\times X});$
$\Phi\mapsto\bigoplus_{\gamma\in\Gamma}(\psi^{\prime}_{\gamma}(\Phi)-(\operatorname{id}_{\mathfrak{C}}\times\gamma)^{*}\Phi\cdot(1,\gamma)^{-1})$
is 00 — note that $\psi^{\prime}_{\gamma}(\Phi)\in\mathcal{H}om((p_{\mathfrak{C}}\times\gamma)^{*}\mathfrak{F}_{a,b,c,\mathfrak{J}},\mathcal{O}_{\mathfrak{C}\times X})$ because we are working on $\mathfrak{C}_{2}$.
Finally, let $\mathfrak{C}_{6}\subset\mathfrak{C}_{1}$ be the subset of points $t$ such that the image of $\psi^{\prime}\oplus\psi^{\prime\prime}$ by the homomorphism
$(p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}(V_{\Gamma,\mathfrak{I}}\oplus V_{\Gamma,\mathfrak{I}}^{\vee})\otimes\mathcal{O}_{\mathfrak{C}}(1)\to V_{\Gamma,\mathfrak{I}}\otimes\mathcal{O}_{\mathfrak{C}}(1);\,\psi^{\prime}\oplus\psi^{\prime\prime}\mapsto\psi^{\prime\prime}-\psi^{\prime*-1},$
where
(27) $\psi^{\prime*-1}=\bigoplus_{\gamma\in\Gamma}\psi^{\prime*-1}_{\gamma},$
is 00 — note that we are using the notation $\psi^{\prime}=\bigoplus_{\gamma\in\Gamma}\psi^{\prime}_{\gamma}$. By [21, Proposition 2.3.5.1], this is a closed subscheme of $\mathfrak{C}_{1}$, so it is locally closed in $\mathfrak{C}$.
Remark 5.4.
To obtain the parameter space of twisted equivariant $G$-bundle, we restrict below to a closed subvariety $\mathfrak{C}_{0}\subset\mathfrak{C}_{6}$. From this point of view we can now see one of the reasons why we work with $\mathbb{P}(V_{\Gamma,\mathfrak{I}}\oplus V^{\vee}_{\Gamma})$, rather than $\mathbb{P}(V_{\Gamma,\mathfrak{I}})$. Indeed, a point in $\mathbb{P}(V_{\Gamma,\mathfrak{I}})$ only determines a set of homomorphisms $V\otimes\mathcal{O}_{X}\to\gamma^{*}V\otimes\mathcal{O}_{X}$ up to rescaling by an element of $\mathbb{C}^{*}$ — in other words, up to a choice of local trivialization of $\mathcal{O}_{\mathfrak{C}}(1)$. However, given a point in $\mathfrak{C}_{6}\subset\mathbb{P}(V_{\Gamma,\mathfrak{I}}\oplus V^{\vee}_{\Gamma})$, these homomorphisms are uniquely defined by the condition that $\psi^{\prime\prime}$ is the “dual inverse” of $\psi^{\prime}$.
Note that every vector bundle $E$ of rank $n$ and trivial determinant with a reduction of structure group to $G$ making it a semistable $G$-bundle, equipped with a set of homomorphisms $E\to\gamma^{*}E$ for $\gamma\in\Gamma$, is represented by a point in $\mathfrak{C}$, since every such homomorphism lifts to a homomorphism $V\to\gamma^{*}V$ via the isomorphism $V\cong H^{0}(E(m))$ between the spaces of global sections. More precisely, any such object is represented by a point in $\mathfrak{C}_{2}$. This implies, by Proposition 2.5, that every semistable pseudo-equivariant $G$-bundle is represented by a point in $\mathfrak{C}_{2}$. Conversely, every point in $\mathfrak{C}_{2}$ represents a — not necessarily semistable — vector bundle $E$ of rank $n$ and trivial determinant with a reduction of structure group to $G$, equipped with a set of homomorphisms $E\to\gamma^{*}E$ for $\gamma\in\Gamma$. Inside $\mathfrak{C}_{1}\cap\mathfrak{C}_{2}$, all the homomorphisms $E\to\gamma^{*}E$ are isomorphisms. Moreover, points in $\mathfrak{C}_{1}\cap\mathfrak{C}_{2}\cap\mathfrak{C}_{3}\cap\mathfrak{C}_{4}$ and $\mathfrak{C}_{1}\cap\mathfrak{C}_{2}\cap\mathfrak{C}_{5}$ satisfy (7) and (8), respectively. Therefore, every point in $\mathfrak{C}_{1}\cap\mathfrak{C}_{2}\cap\mathfrak{C}_{3}\cap\mathfrak{C}_{4}\cap\mathfrak{C}_{5}$ represents a pseudo-equivariant $G$-bundle, and any semistable pseudo-equivariant $G$-bundle is represented by a point in $\mathfrak{C}_{1}\cap\mathfrak{C}_{2}\cap\mathfrak{C}_{3}\cap\mathfrak{C}_{4}\cap\mathfrak{C}_{5}$. Via Proposition 3.3, the same statement is true after replacing pseudo-equivariant $G$-bundles with twisted equivariant $G$-bundles. Moreover, if we restrict to $\mathfrak{C}_{1}\cap\mathfrak{C}_{2}\cap\mathfrak{C}_{3}\cap\mathfrak{C}_{4}\cap\mathfrak{C}_{5}\cap\mathfrak{C}_{6}$, then all the semistable twisted equivariant $G$-bundles are still represented, and each set of isomorphisms $E\to\gamma^{*}E$ only appears once for each $G$-bundle. Set $\mathfrak{C}_{0}:=\mathfrak{C}_{1}\cap\mathfrak{C}_{2}\cap\mathfrak{C}_{3}\cap\mathfrak{C}_{4}\cap\mathfrak{C}_{5}\cap\mathfrak{C}_{6}$.
Let $\mathfrak{F}_{\mathfrak{C}}:=(p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}\mathfrak{F}_{\mathfrak{I}}$, $\mathfrak{F}_{\mathfrak{C}_{0}}:=(p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}\mathfrak{F}_{\mathfrak{I}}|_{\mathfrak{C}_{0}\times X}$ and $\varphi_{\mathfrak{C}_{0}}:=(p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}\varphi|_{\mathfrak{C}_{0}\times X}\in\operatorname{Hom}((\mathfrak{F}_{\mathfrak{C}_{0}})_{a,b,c},\mathcal{O}_{\mathfrak{C}\times X})$. Denote by
$\psi_{\mathfrak{C}_{0}}\in H^{0}\left(\bigoplus_{\gamma\in\Gamma}\mathcal{H}om(\mathfrak{F}_{\mathfrak{C}_{0}},(\operatorname{id}_{\mathfrak{C}_{0}}\times\gamma)^{*}\mathfrak{F}_{\mathfrak{C}_{0}})\right)$
the section induced by $\psi^{\prime}$. Then $(\mathfrak{F}_{\mathfrak{C}_{0}},\varphi_{\mathfrak{C}_{0}},\psi_{\mathfrak{C}_{0}})$ is a universal family of pseudo-equivariant pairs over $\mathfrak{C}_{0}\times X$. Let $V_{\mathfrak{C}}:=V\otimes\mathcal{O}_{\mathfrak{C}\times X}$, which is equipped with a surjective homomorphism $q_{\mathfrak{C}}:=(p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}q_{\mathfrak{I}}:V_{\mathfrak{C}}(-m)\to\mathfrak{F}_{\mathfrak{C}}$.
The next proposition reduces the problem of constructing the moduli space of twisted equivariant $G$-bundles to the study of a GIT quotient $\mathfrak{C}_{0}\sslash\mathrm{GL}(V)$.
Proposition 5.5.
The natural action of $\mathrm{GL}(V)$ on $V$ induces an action on $\mathfrak{C}$ preserving $\mathfrak{C}_{0}\subset\mathfrak{C}$. For every $t$ and $s\in\mathfrak{C}_{0}$, the pseudo-equivariant $\rho_{a,b,c}$-pairs $(\mathfrak{F}_{\mathfrak{C}_{0}},\varphi_{\mathfrak{C}_{0}},\psi_{\mathfrak{C}_{0}})|_{\{t\}\times X}$ and $(\mathfrak{F}_{\mathfrak{C}_{0}},\varphi_{\mathfrak{C}_{0}},\psi_{\mathfrak{C}_{0}})|_{\{s\}\times X}$ are isomorphic if and only if they are in the same $\mathrm{GL}(V)$-orbit.
Remark 5.6.
By Propositions 3.3 and 4.9, two twisted equivariant $G$-bundles are isomorphic if and only if the corresponding pseudo-equivariant pairs are, so Proposition 5.5 implies that two twisted equivariant $G$-bundles are isomorphic if and only if the corresponding points of $\mathfrak{C}_{0}$ are in the same $\mathrm{GL}(V)$-orbit.
Proof.
The $\mathrm{GL}(V)$-invariance of $\mathfrak{C}_{0}$ follows from its definition. Given $t\in\mathfrak{C}_{0}$, consider the quotient $q_{\mathfrak{C}}|_{\{t\}\times X}:V_{\mathfrak{C}}(-m)|_{\{t\}\times X}\to\mathfrak{F}_{\mathfrak{C}}|_{\{t\}\times X}$. An element $g\in\mathrm{GL}(V)$ sends $\ker q_{\mathfrak{C}}|_{\{t\}\times X}\subset V(-m)$, which is a subbundle of rank $\dim V-n$, to another subbundle of $V(-m)$, and the quotient is equal to $q_{\mathfrak{C}}|_{\{s\}\times X}:V_{\mathfrak{C}}(-m)|_{\{s\}\times X}\to\mathfrak{F}_{\mathfrak{C}}|_{\{s\}\times X}$ for some $s\in\mathfrak{C}$. Moreover, the automorphism of $V_{\mathfrak{C}}$ determined by $g$ induces an isomorphism of vector bundles $\mathfrak{F}_{\mathfrak{C}_{0}}|_{\{t\}\times X}\xrightarrow{\sim}\mathfrak{F}_{\mathfrak{C}_{0}}|_{\{s\}\times X}$. Via this isomorphism, the pseudo-equivariant pair $(\mathfrak{F}_{\mathfrak{C}_{0}},\varphi_{\mathfrak{C}_{0}},\psi_{\mathfrak{C}_{0}})|_{\{t\}\times X}$ determines another one $(\mathfrak{F}_{\mathfrak{C}_{0}},\varphi_{\mathfrak{C}_{0}},\psi_{\mathfrak{C}_{0}})|_{\{s\}\times X}$, which determines $s$ uniquely.
It is left to prove that, given $t$ and $s\in\mathfrak{C}$ such that the corresponding pseudo-equivariant pairs $(\mathfrak{F}_{\mathfrak{C}_{0}},\varphi_{\mathfrak{C}_{0}},\psi_{\mathfrak{C}_{0}})|_{\{t\}\times X}$ and $(\mathfrak{F}_{\mathfrak{C}_{0}},\varphi_{\mathfrak{C}_{0}},\psi_{\mathfrak{C}_{0}})|_{\{s\}\times X}$ are isomorphic, they are in the same $\mathrm{GL}(V)$-orbit. An isomorphism induces an isomorphism of vector spaces $H^{0}(\mathfrak{F}_{\mathfrak{C}}(m)|_{\{t\}\times X})\cong H^{0}(\mathfrak{F}_{\mathfrak{C}}(m)|_{\{s\}\times X})$ which in turn, via the isomorphism of vector spaces of sections $H^{0}(q_{\mathfrak{C}}\otimes\operatorname{id}_{\mathcal{O}_{X}(m)})$, induces an automorphism of $V$, i.e. an element of $\mathrm{GL}(V)$. It is clear that this induces the isomorphism that we started with, as required. ∎
6. The moduli space for $G$ semisimple
We maintain the notation of the previous section. In order to perform a GIT quotient $\mathfrak{C}_{0}\sslash\mathrm{GL}(V)$ providing a moduli space parametrizing isomorphism classes of polystable $(\theta,c)$-twisted equivariant $G$-bundles, we need to find a linearization of the $\mathrm{GL}(V)$-action on some very ample line bundle over $\mathfrak{C}_{0}$, such that the corresponding GIT definition of (semi, poly)stablity is equivalent to the definitions of Section 2.1. This section is devoted to showing that such a linearization exists, and that the resulting moduli space is a complex projective variety.
6.1. Gieseker space and linearization
In order to find an appropriate linearization of the $\mathrm{GL}(V)$-action on $\mathfrak{C}_{0}$ and to apply the Hilbert–Mumford criterion, we use an injective morphism into a projective space, usually called a Gieseker space — the role of Gieseker spaces in the construction of the moduli space of $G$-bundles is explained in [21, Section 2.3.5].
Let $\mathcal{G}{(m)}_{1}:=\operatorname{Hom}(\bigwedge^{n}V\otimes\mathcal{O}_{X},\mathcal{O}_{X}(nm))$ and $\mathbb{G}{(m)}_{1}:=\mathbb{P}(\mathcal{G}{(m)}_{1})$ — a complex projective space, equipped with the very ample line bundle $\mathcal{O}_{\mathbb{G}{(m)}_{1}}(1)$. With notation as in Section 5.3, let $\mathcal{L}_{\mathfrak{C}}\to\mathfrak{C}$ be the line bundle that satisfies $\bigwedge^{n}\mathfrak{F}_{\mathfrak{C}}\cong\pi_{\mathfrak{C}}^{*}\mathcal{L}_{\mathfrak{C}}$ — this exists because the determinant of $\mathfrak{F}_{\mathfrak{C}}|_{\{s\}\times X}$ is trivial for every $s\in\mathfrak{C}$, see Section 5.1. Then the surjective homomorphism $q_{\mathfrak{C}}:V_{\mathfrak{C}}(-m)\to\mathfrak{F}_{\mathfrak{C}}$ induces a homomorphism
$\bigwedge^{n}(q_{\mathfrak{C}}\otimes\operatorname{id}_{\pi_{X}^{*}\mathcal{O}_{X}(nm)}):\bigwedge^{n}V\otimes\mathcal{O}_{\mathfrak{C}\times X}\to\pi_{X}^{*}\mathcal{O}_{X}(nm)\otimes\pi_{\mathfrak{C}}^{*}\mathcal{L}_{\mathfrak{C}}.$
This, in turn, determines a morphism
$\operatorname{Gies}_{1}:\mathfrak{C}\to\mathbb{G}{(m)}_{1}$
such that $\operatorname{Gies}_{1}^{*}\mathcal{O}_{\mathbb{G}{(m)}_{1}}(1)\cong\mathcal{L}_{\mathfrak{C}}$.
Now let $V_{a,b}:=\left(V^{\otimes a}\right)^{\oplus b}=V_{a,b,0}$,
$\mathcal{G}{(m)}_{2}:=\operatorname{Hom}\left(V_{a,b}\otimes\mathcal{O}_{X},\mathcal{O}_{X}(ma)\right)$
and $\mathbb{G}{(m)}_{2}:=\mathbb{P}(\mathcal{G}{(m)}_{2})$ — a complex projective space, equipped with the very ample line bundle $\mathcal{O}_{\mathbb{G}{(m)}_{2}}(1)$. Then the homomorphism $\varphi_{\mathfrak{C}}:(\mathfrak{F}_{\mathfrak{C}})_{a,b,c}\to\pi_{X}^{*}\mathcal{O}_{X}$ induces a homomorphism
$V_{a,b}\otimes\mathcal{O}_{\mathfrak{C}\times X}\to(\mathfrak{F}_{\mathfrak{C}})_{a,b}\otimes\pi_{X}^{*}(\mathcal{O}_{X}(ma))\to\pi_{X}^{*}(\mathcal{O}_{X}(ma))\otimes\pi_{\mathfrak{C}}^{*}(\mathcal{L}^{\otimes c}_{\mathfrak{C}}),$
which in turn induces a morphism $\operatorname{Gies}_{2}:\mathfrak{C}\to\mathbb{G}{(m)}_{2}$ such that $\operatorname{Gies}_{2}^{*}\mathcal{O}_{\mathbb{G}{(m)}_{2}}(1)\cong\mathcal{L}_{\mathfrak{C}}^{\otimes c}$.
Finally, let
$V_{\Gamma}:=\bigoplus_{\gamma\in\Gamma}\operatorname{Hom}(V\otimes\mathcal{O}_{X},\gamma^{*}(V\otimes\mathcal{O}_{X}))\quad\text{and}\quad V_{\Gamma}^{\vee}:=\bigoplus_{\gamma\in\Gamma}\operatorname{Hom}(V^{*}\otimes\mathcal{O}_{X},\gamma^{*}(V^{*}\otimes\mathcal{O}_{X})),$
and set $\mathcal{G}{(m)}_{3}:=V_{\Gamma}\oplus V^{\vee}_{\Gamma}$ and $\mathbb{G}{(m)}_{3}:=\mathbb{P}(\mathcal{G}{(m)}_{3})$. The tautological section $\psi^{\prime}\oplus\psi^{\prime\prime}\in(p_{\mathfrak{C}}\times\operatorname{id}_{X})^{*}(V_{\Gamma,\mathfrak{I}}\oplus V_{\Gamma,\mathfrak{I}}^{\vee})$ determines a morphism $\operatorname{Gies}_{3}:\mathfrak{C}\to\mathbb{G}{(m)}_{3}$.
The product
$\operatorname{Gies}:=\operatorname{Gies}_{1}\times\operatorname{Gies}_{2}\times\operatorname{Gies}_{3}:\mathfrak{C}\to\mathbb{G}{(m)}:=\mathbb{G}{(m)}_{1}\times\mathbb{G}{(m)}_{2}\times\mathbb{G}{(m)}_{3}$
is an injective $\mathrm{GL}(V)$-equivariant morphism. Of course, it restricts to an injective morphism on the locally closed subvariety $\mathfrak{C}_{0}$. We refer to $\mathbb{G}{(m)}$ as the Gieseker space, and to $\operatorname{Gies}$ as the Gieseker morphism.
Given positive integers $s_{1},s_{2}$ and $\chi$, let $\mathcal{O}_{\mathbb{G}{(m)}}(s_{1},s_{2},s_{1}\chi):=\mathcal{O}_{\mathbb{G}{(m)}_{1}}(s_{1})\otimes\mathcal{O}_{\mathbb{G}{(m)}_{2}}(s_{2})\otimes\mathcal{O}_{\mathbb{G}{(m)}_{3}}(s_{1}\chi)$, a very ample line bundle. For each triple $(s_{1},s_{2},\chi)$ the natural $\mathrm{GL}(V)$-action on $\mathcal{O}_{\mathbb{G}{(m)}}(s_{1},s_{2},s_{1}\chi)$ determines a linearization of the $\mathrm{GL}(V)$-action on $\mathbb{G}{(m)}$ — and, therefore, a linearization of the $\mathrm{GL}(V)$-action on $\mathfrak{C}$.
Note that, for each 1-parameter subgroup $\lambda:\mathbb{C}^{*}\to\mathrm{GL}(V)$ and every point $t=(t_{1},t_{2},t_{3})\in\mathbb{G}{(m)}$,
$\mu(t,\lambda)=s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)+s_{1}\chi\mu(t_{3},\lambda),$
where $\mu$ is defined by taking representatives of $t,t_{1},t_{2}$ and $t_{3}$ in the respective vector spaces and using (12). Note also that, by Remark 4.3, if $t_{3}=[\Psi^{\prime},\Psi^{\prime\prime}]\in\mathbb{G}{(m)}_{3}$ for some $\Psi^{\prime}\in V_{\Gamma}$ and $\Psi^{\prime\prime}\in V^{\vee}_{\Gamma}$, then $\mu(t_{3},\lambda)=\max\left(\mu(\Psi^{\prime},\lambda),\mu(\Psi^{\prime\prime},\lambda)\right)$. If we further assume that $t\in\operatorname{Gies}(\mathfrak{C}_{0})$, then $\mu(\Psi^{\prime\prime},\lambda)=\mu(\Psi^{\prime*-1},\lambda)$ and we may argue that $\mu(t_{3},\lambda)\geq 0$ as in the proof of Lemma 4.12, with equality if and only if $\lambda$ is $\Gamma$-invariant with respect to the $\Gamma$-equivariant structure on $V\otimes\mathcal{O}_{X}$ determined by $\Psi^{\prime}$.
Given $\delta>0$, set
(28) $\epsilon:=\frac{s_{1}}{s_{2}}=\frac{\dim V-a\delta}{n\delta}.$
Similarly to the proof of Lemma 4.6, there exists a finite set $S$ of 1-parameter subgroups $\lambda:\mathbb{C}^{*}\to\mathrm{GL}(V)$ such that $t=(t_{1},t_{2},t_{3})\in\mathbb{G}{(m)}$ if (semi)stable if and only if $\mu(g\cdot t,\lambda)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$ for every $g\in\mathrm{GL}(V)$ and $\lambda\in S$. We may further assume that all the elements in $S$ have integral weights.
Since the weights of each $\lambda\in S$ are bounded, given a fixed $\lambda\in S$, $\mu(g\cdot t_{i},\lambda)$ is bounded for each $\lambda\in S$ and $i=1,2,3$. In particular, we may choose $\chi\gg\epsilon^{-1}\gg 0$ so that, for every 1-parameter subgroup in $S$ and every point $t\in\mathbb{G}{(m)}$, $\mu(t,\lambda)>0$ if $\mu(t_{3},\lambda)>0$ for each $\lambda\in S$. Therefore, if $t\in\operatorname{Gies}(\mathfrak{C}_{0})$ and $\lambda\in S$, then $\mu(t,\lambda)>0$ if and only if either $\mu(t_{3},\lambda)>0$, or $\mu(t_{3},\lambda)=0$ and $s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)>0$, with $\mu(t_{3},\lambda)=0$ if and only if $\lambda$ is $\Gamma$-invariant with respect to the $\Gamma$-equivariant structure defined by $t_{3}$. In particular, $\mu(t,\lambda)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$ for every 1-parameter $\lambda$ if and only if $\mu(t,\lambda)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$ for every $\Gamma$-invariant $\lambda$.
With the above linearization, the Gieseker map allows us to study the GIT quotient $\mathfrak{C}_{0}\sslash\mathrm{GL}(V)$ by means of the GIT quotient $\mathbb{G}{(m)}\sslash\mathrm{GL}(V)$, where $\mathbb{G}{(m)}$ is a projective space.
6.2. Hilbert–Mumford criterion vs (semi)stability
Denote by $\mathfrak{G}(m_{1})$ the set of isomorphism classes of vector bundles $E$ of rank $n$ and trivial determinant for which there exist $m\geq m_{1}$ and $t\in\mathfrak{C}_{0}$ such that $E\cong\mathfrak{F}_{\mathfrak{C}}|_{\{t\}\times X}$ and $\operatorname{Gies}(t)$ is semistable with respect to the linearization in Section 6.1. In what follows we assume that $\chi\gg\epsilon^{-1}\gg 0$ or, alternatively, $\chi\gg\delta\gg 0$ — see Proposition 4.11 and the end of Section 6.1 on why we need these assumptions.
Lemma 6.1.
There is $m_{1}>0$ and $C\in\mathbb{Q}$, such that the following holds true: For every $E\in\mathfrak{G}(m_{1})$, and every $\Gamma$-invariant vector subbundle $F\subset E$ — with respect to the $\Gamma$-equivariant structure given by a point $t\in\mathfrak{C}_{0}$ corresponding to $E$ —, $\mu(F)\leq C$.
Proof.
Fix a vector bundle $E\in\mathfrak{G}$ with corresponding point $t\in\mathfrak{C}_{0}$. Following the proof of [21, 2.3.5.12], we may find a lower bound for
[…]
Indeed, fix a $\Gamma$-invariant subbundle $F\subset E$. There is an exact sequence
$0\to H^{0}(F(m))\to H^{0}(E(m))\to H^{0}(Q(m)),$
where $Q:=E/F$. Take a 1-parameter subgroup $\lambda:\mathbb{C}^{*}\to\mathrm{SL}(V)$ whose corresponding weighted filtration is
$\left(V_{\bullet}(\lambda):\{0\}\subset V_{1}:=H^{0}(q(m))^{-1}\left(H^{0}(F(m))\right)\subset V,\;\alpha_{\bullet}(\lambda)=(1)\right).$
Since $F$ is $\Gamma$-invariant, so is $V_{1}$, the weighted filtration $V_{\bullet}(\lambda)$ and $\lambda$ itself. Hence, if $\operatorname{Gies}(t)=(t_{1},t_{2},t_{3})$, then $\mu(t_{3},\lambda)=0$. In particular, $\mu(\operatorname{Gies}(t),\lambda)=s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)$. Since $\operatorname{Gies}(t)$ is semistable, $s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)\geq 0$. Applying the same calculation as in the proof of [21, 2.3.5.12], we conclude that $\mu_{\text{min}}^{\Gamma}(E)\geq-\frac{\delta a}{n}-g$, as required. ∎
Lemma 6.2.
There exists a positive integer $m_{2}\geq m_{1}$ and a finite set
$S\subset\{(n_{\bullet},\alpha_{\bullet})\;:\;n_{j}\in\mathbb{N},\alpha_{j}\in\mathbb{Q}_{\geq 0},\sum_{j}n_{j}=n\},$
such that any $m\geq m_{2}$ satisfies the following. For every pseudo-equivariant pair $(E,\sigma,\varphi)$ such that $E\in\mathfrak{G}(m_{2})$, the following conditions are equivalent.
  1. (1)
    $(E,\sigma,\varphi)$ is $(\delta,\chi)$-(semi)stable.
  2. (2)
    For every $\Gamma$-invariant weighted filtration $(E_{\bullet},\alpha_{\bullet})$ with $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$ such that $E_{j}(m)$ is globally generated and $h^{1}(E_{j}(m))=0$ for each $i$,
    $\sum_{j}\alpha_{j}\left(h^{0}(E(m))\operatorname{rk}(E_{j})-h^{0}(E_{j}(n))\operatorname{rk}(E)\right)+\delta\mu(E_{\bullet},\alpha_{\bullet},\sigma)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0.$
Proof.
By Lemma 4.6, there is a finite set
$S\subset\{(n_{\bullet},\alpha_{\bullet})\;:\;n_{j}\in\mathbb{N},\alpha_{j}\in\mathbb{Q}_{\geq 0},\sum_{j}n_{j}=n\}$
such that the $(\delta,\chi)$-(semi)stability of any pseudo-equivariant pair $(E,\sigma,f)$ only has to be verified for weighted filtrations $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$. Since $\chi\gg\delta\gg 0$, this only has to be verified for $\Gamma$-invariant weighted filtrations $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$. Fix a constant $C^{\prime}$. Then, there exists $C^{\prime\prime}$ such that, for every pseudo-equivariant pair $(E,\sigma,f)$ with $E\in\mathfrak{G}(m_{1})$ and every $\Gamma$-invariant weighted filtration $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$, if
(29) $\mu(E_{j})\leq C^{\prime\prime}$
for some $i$, then $M(E_{\bullet},\alpha_{\bullet})>C^{\prime}$. This follows directly from $\mu(E_{j})\leq\mu_{\text{max}}^{\Gamma}(E)$ and the uniform bound on $\mu_{\text{max}}^{\Gamma}(E)$ given by Lemma 6.1, where $\mu_{\text{max}}^{\Gamma}(E)$ is the maximum slope attained by a $\Gamma$-invariant vector subbundle of $E$. Using [21, Lemma 1.5.1.41], we may determine a constant $C^{\prime\prime\prime}$ such that $\mu(E_{\bullet},\alpha_{\bullet},\sigma)\geq-C^{\prime\prime\prime}$ for every $\Gamma$-invariant weighted filtration $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$. If we choose $C^{\prime}\geq\delta C^{\prime\prime\prime}$, then
$M(E_{\bullet},\alpha_{\bullet})+\delta\mu(E_{\bullet},\alpha_{\bullet},\sigma)>C^{\prime}-\delta C^{\prime\prime\prime}\geq 0.$
Therefore, $(\delta,\chi)$-(semi)stability only has to be verified for $\Gamma$-invariant weighted filtrations $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$ and (29) fails. By a slightly refined version of [21, Proposition 2.2.3.7], these weighted filtrations live in bounded families, hence the lemma follows by [21, Proposition 2.2.3.2]. ∎
Proposition 6.3.
There is $m_{3}\geq m_{2}$ such that, for every $m\geq m_{3}$ and every point $t\in\mathfrak{C}_{0}$ with (semi)stable Gieseker point $\operatorname{Gies}(t)\in\mathbb{G}{(m)}$, the corresponding pseudo-equivariant pair $(E_{t},\sigma_{t},f_{t})$ is $(\delta,\chi)$-semistable.
Proof.
We just need to check condition (2) in the statement of Lemma 6.2, applied to $(E_{t},\sigma_{t},f_{t})$. To do so, follow the same argument as in the proof of [21, Theorem 2.3.5.13] after the statement of [21, Corollary 2.3.5.14]. ∎
Recall that, by construction, for every $(\delta,\chi)$-(semi)stable pseudo-equivariant pair $(E,\sigma,f)$ corresponding to a twisted equivariant $G$-bundle via Propositions 3.3 and 4.9, there exists $t\in\mathfrak{C}_{0}$ such that $(E,\sigma,f)\cong(\mathfrak{F}_{\mathfrak{C}_{0}}|_{\{t\}\times X},\varphi_{\mathfrak{C}_{0}}|_{\{t\}\times X},\psi_{\mathfrak{C}_{0}}|_{\{t\}\times X})$ — see Section 5.3.
Lemma 6.4.
There exists a positive integer $m_{4}$, such that any $(\delta,\chi)$-(semi)stable pseudo-equivariant pair $(E,\sigma,f)$ satisfies
$\sum_{j}\alpha_{j}(\dim V\cdot\operatorname{rk}E_{j}-h^{0}(E_{j}(m))\cdot n)+\delta\cdot\mu(E_{\bullet},\alpha_{\bullet},\sigma)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$
for every $\Gamma$-invariant weighted filtration $(E_{\bullet},\alpha_{\bullet})$ of $E$ and every $m\geq m_{4}$
Proof.
By the proof of [21, Proposition 2.3.5.15], the lemma follows if we can prove that $M(E_{\bullet}^{A},\alpha_{\bullet}^{A})+\delta\mu(E_{\bullet}^{A},\alpha_{\bullet}^{A},\sigma)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$ for an arbitrary $\Gamma$-invariant weighted filtration $(E_{\bullet},\alpha_{\bullet})$ — indeed, the last inequality of the last computation in the proof is the only place where the (semi)stability assumptions are used. Note that $\Gamma$-invariance of $(E_{\bullet},\alpha_{\bullet})$ implies, in particular, that $(E_{\bullet}^{A},\alpha_{\bullet}^{A})$ is $\Gamma$-invariant. By Lemma 4.12 and Definition 4.4, the $(\delta,\chi)$-(semi)stability of $(E,\sigma,f)$ implies that $M(E_{\bullet}^{A},\alpha_{\bullet}^{A})+\delta\mu(E_{\bullet}^{A},\alpha_{\bullet}^{A},\sigma)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$, as required. ∎
Proposition 6.5.
There exists a positive integer $m_{5}$ satisfying the following property. If $m\geq m_{5}$ and $(E,\sigma,f)$ is a $(\delta,\chi)$-(semi)stable pseudo-equivariant pair then, given $t\in\mathfrak{C}_{0}$ such that $(E,\sigma,f)\cong(\mathfrak{F}_{\mathfrak{C}}|_{\{t\}\times X},\varphi_{\mathfrak{C}}|_{\{t\}\times X},\psi_{\mathfrak{C}}|_{\{t\}\times X})$, the associated Gieseker point $\operatorname{Gies}(t)$ is (semi)stable for the linearization given in Section 6.1.
Proof.
Since $\chi\gg\epsilon^{-1}\gg 0$, it is enough to check that $s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2}),\lambda)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$ for all non-trivial $\Gamma$-invariant 1-parameter subgroups $\lambda$. To do this, just copy the proof of [21, Theorem 2.3.5.16] and replace the arbitrary $\lambda$ with a $\Gamma$-invariant $\lambda$, and [21, Proposition 2.3.5.15] by Lemma 6.4. This works because the $\Gamma$-invariance of $\lambda$ clearly implies that the associated filtration $E_{\bullet}$ is $\Gamma$-invariant. ∎
Theorem 6.6.
There exists a positive integer $m_{0}$ satisfying the following property. If $m\geq m_{0}$ and $\chi\gg\delta\gg 0$, then, given $t\in\mathfrak{C}_{0}$ and $(E,\sigma,f):=(\mathfrak{F}_{\mathfrak{C}}|_{\{t\}\times X},\varphi_{\mathfrak{C}}|_{\{t\}\times X},\psi_{\mathfrak{C}}|_{\{t\}\times X})$, the pseudo-equivariant pair $(E,\sigma,f)$ is $(\delta,\chi)$-(semi)stable if and only if $\operatorname{Gies}(t)$ is (semi)stable for the linearization given in Section 6.1.
Corollary 6.7.
There exists a complex quasi-projective variety $\mathfrak{C}_{0}^{ss}$, together with a universal $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundle $(\mathcal{Q},\bullet)$ over $\mathfrak{C}_{0}^{ss}\times X$ such that, for every semistable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundle $(Q,\bullet)$ over $X$, there exists $t\in\mathfrak{C}_{0}^{ss}$ with $(Q,\bullet)\cong(\mathcal{Q}|_{\{t\}\times X},\bullet)$. Moreover, $\mathfrak{C}_{0}^{ss}$ features an algebraic $\mathrm{GL}(V)$-action lifting to a $\mathrm{GL}(V)$-action on $\mathcal{Q}$, such that $(\mathcal{Q}|_{\{t\}\times X},\bullet)\cong(\mathcal{Q}|_{\{t^{\prime}\}\times X},\bullet)$ if and only if $t$ and $t^{\prime}$ are in the same $\mathrm{GL}(V)$-orbit.
Proof.
By Theorem 6.6, the subset $\mathfrak{C}_{0}^{ss}\subset\mathfrak{C}_{0}$ consisting of $(\delta,\chi)$-semistable pseudo-equivariant pairs is an open subvariety. The result follows from Propositions 5.5, 4.11 and 3.3. ∎
6.3. The moduli space is a projective variety
Let $\mathbb{G}{(m)}^{ss}\subset\mathbb{G}{(m)}$ be the open subvariety of semistable points in $\mathbb{G}{(m)}$ in the sense of GIT, with respect to the $\mathrm{GL}(V)$-action and the linearization given in Section 6.1. Choose $\chi\gg\delta\gg 0$ and $\chi\gg\epsilon^{-1}\gg 0$ so that Propositions 6.3 and 6.5 hold. Let $\mathfrak{C}_{0}^{ss}\subset\mathfrak{C}_{0}$ be the subvariety consisting of points corresponding to $(\delta,\chi)$-semistable pseudo-equivariant pairs. Pick $m_{0}$ as in Theorem 6.6. Then, for $m\geq m_{0}$, the Gieseker morphism restricts to an injection $\operatorname{Gies}^{ss}:\mathfrak{C}_{0}^{ss}\hookrightarrow\mathbb{G}{(m)}^{ss}$.
Proposition 6.8.
The Gieseker morphism $\operatorname{Gies}^{ss}:\mathfrak{C}_{0}^{ss}\hookrightarrow\mathbb{G}{(m)}^{ss}$ is proper.
Proof.
We use the valuative criterion for properness. Let $R$ be a discrete valuation ring with quotient field $K$ and set $C:=\operatorname{Spec}(R)$ and $C^{*}:=\operatorname{Spec}(K)$. Let $\eta:C\to\mathbb{G}{(m)}^{ss}$ be a morphism such that the restriction $\eta|_{C^{*}}:C^{*}\to\mathbb{G}{(m)}^{ss}$ lifts to a morphism $\eta^{*}:C^{*}\to\mathfrak{C}_{0}^{ss}$. We have to show that $\eta^{*}$ extends to a morphism $C\to\mathfrak{C}_{0}^{ss}$ lifting $\eta$.
Since $\mathfrak{C}_{0}^{ss}$ is a parameter space for the set of $(\delta,\chi)$-semistable pseudo-equivariant pairs — see Section 5.3 —, the morphism $\eta^{*}$ is associated to a family of $(\delta,\chi)$-semistable pseudo-equivariant pairs
$(q_{C^{*}}:V\otimes\pi_{X}^{*}(\mathcal{O}_{X}(-m))\otimes\pi_{C^{*}}^{*}\mathcal{O}_{C^{*}}\to E_{C^{*}},\sigma_{C^{*}},f_{C^{*}}).$
Here $E_{C^{*}}\to C^{*}\times X$ is a family of vector bundles of rank $n$ and trivial determinant, $\sigma_{C^{*}}$ is a homomorphism
$\sigma_{C^{*}}:(E_{C^{*}})_{a,b}=(E^{\otimes a})^{\oplus b}\to(\det E)^{\otimes c}\otimes\pi_{X}^{*}\mathcal{O}_{X},$
and $(\operatorname{id}_{C^{*}}\times\gamma)^{*}E_{C^{*}}\}_{\gamma\in\Gamma}$ is a family of isomorphisms satisfying 7 and 11.
Recall that the Quot Scheme $\mathfrak{Q}$ has a natural compactification $\overline{\mathfrak{Q}}$ parametrizing coherent sheaves of rank $n$. Therefore, $q_{C^{*}}$ extends to a family $q_{C}:V\otimes\pi_{X}^{*}(\mathcal{O}_{X}(-m))\to E_{C}$, where the restriction $E_{C}|_{\{0\}\times X}$ to the special fibre may have torsion at a subset of points $Z\subset X$. However, note that $\widetilde{E}_{C}:=E_{C}^{**}$, is a family of locally-free sheaves, since $\widetilde{E}_{C}$ is a reflexive sheaf over a the regular two-dimensional scheme $C\times X$ — see [16, Corollaries 1.3 and 1.4].
By the proof of [21, Proposition 2.3.5.17], $\sigma_{C^{*}}$ extends to a family of homomorphisms $\sigma_{C}:(E_{C})_{a,b,c}\to\pi_{X}^{*}\mathcal{O}_{X}$. As in such proof, let $\iota:(C\times X)\setminus(C\times Z)\hookrightarrow C\times X$ be the open embedding and let
$\widetilde{\sigma}_{C}:(\widetilde{E}_{C})_{a,b}\to\iota_{*}\left((\widetilde{E}_{C})_{a,b}|_{(C\times X)\setminus(C\times Z)}\right)=\iota_{*}\left((E_{C})_{a,b}|_{(C\times X)\setminus(C\times Z)}\right)$
$\xrightarrow{\iota_{*}(\sigma_{C}|_{(C\times X)\setminus(C\times Z)})}\iota_{*}\left(\det(E_{C}|_{(C\times X)\setminus(C\times Z)})^{\otimes c}\otimes\pi_{X}^{*}\mathcal{O}_{X}|_{(C\times X)\setminus(C\times Z)}\right)=\det(\widetilde{E}_{C})^{\otimes c}\otimes\pi_{X}^{*}\mathcal{O}_{X}.$
Consider the projection $\eta_{3}:C\to\mathbb{G}{(m)}_{3}=\mathbb{P}(V_{\Gamma}\oplus V^{\vee}_{\Gamma})$ to the third component of $\mathbb{G}{(m)}$. Claim: the point $\eta_{3}(0)\in\mathbb{P}(V_{\Gamma}\oplus V^{\vee}_{\Gamma})$ is equal to the class of $\bigoplus_{\gamma\in\Gamma}f_{0,\gamma}\oplus\bigoplus_{\gamma\in\Gamma}f^{*-1}_{0,\gamma}$ in projective space, where $f_{0}$ is a family of isomorphisms satisfying (7) and (11) — recall the notation (27). Indeed, $\eta_{3}(C^{*})$ is equal to the image of the family $\bigoplus_{\gamma\in\Gamma}f_{C^{*},\gamma}\oplus\bigoplus_{\gamma\in\Gamma}f^{*-1}_{C^{*},\gamma}$ in projective space, which we denote by $\left[\bigoplus_{\gamma\in\Gamma}f_{C^{*},\gamma}\oplus\bigoplus_{\gamma\in\Gamma}f^{*-1}_{C^{*},\gamma}\right]$. This has an extension to a family $\left[\bigoplus_{\gamma\in\Gamma}f_{C,\gamma}\oplus\bigoplus_{\gamma\in\Gamma}f^{*-1}_{C,\gamma}\right]$, where $f_{0,\gamma}$ may not be an isomorphism for some $\gamma$. This, in turn, induces a family $\left[\bigoplus_{\gamma\in\Gamma}\widetilde{f}_{C,\gamma}\oplus\bigoplus_{\gamma\in\Gamma}\widetilde{f}^{*-1}_{C,\gamma}\right]$, where $\widetilde{f}_{C,\gamma}:\widetilde{E}_{C}\to(\operatorname{id}_{C}\times\gamma)^{*}\widetilde{E}_{C}$.
Pick $\gamma\in\Gamma$. Since (7) is a closed condition,
$(\gamma^{*(|\gamma|-1)}\widetilde{f}_{0,\gamma})\circ(\gamma^{*(|\gamma|-2)}\widetilde{f}_{0,\gamma})\circ\dots\circ(\gamma^{*}\widetilde{f}_{0,\gamma})\circ\widetilde{f}_{0,\gamma}=\widetilde{f}_{0,1}=\operatorname{id}.$
In particular, $\widetilde{f}_{0,\gamma}$ is an isomorphism for every $\gamma\in\Gamma$. Since (11) is a closed condition, the claim is proved.
Finally, as in the proof of [21, Proposition 2.3.5.17], one can check that $(\widetilde{E}_{0},\widetilde{\sigma}_{0},\widetilde{f}_{0})$ is a $(\delta,\chi)$-semistable pseudo-equivariant pair. Since its image in $\mathbb{G}{(m)}^{ss}$ is equal to $\eta(0)$, it determines a extension $\eta^{*}:C\to\mathfrak{C}_{0}^{ss}$ lifting $\eta$, as required. ∎
6.4. S-equivalence vs polystability
Recall that the GIT quotient $\mathfrak{C}_{0}^{ss}\sslash\mathrm{GL}(V)$ parametrizes S-equivalence classes of semistable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$. Here, two points of $\mathfrak{C}_{0}^{ss}$ are said to be S-equivalent if the closure of their $\mathrm{GL}(V)$-orbits intersect. In the closure of each orbit there is a unique closed orbit, so we may also see $\mathfrak{C}_{0}^{ss}\sslash\mathrm{GL}(V)$ as the moduli space of isomorphism classes of $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$ whose corresponding $\mathrm{GL}(V)$-orbit is closed. The goal of this section is to show that these are precisely the polystable ones, according to Definition 2.3.
Lemma 6.9.
Assume that $\chi\gg\delta\gg 0$, as usual. Let $t\in\mathfrak{C}_{0}^{ss}$ with corresponding pseudo-equivariant pair $(E_{t},\sigma_{t},f_{t})$. Let $\lambda:\mathbb{C}^{*}\to\mathrm{GL}(V)$ be a non-trivial 1-parameter subgroup with corresponding weighted filtration $(E_{t\bullet},\alpha_{\bullet})$. Then, $\mu(\operatorname{Gies}(t),\lambda)=0$ — according to the linearization given by $\mathcal{O}(s_{1},s_{2},s_{1}\chi)$, see Section 6.1 — if and only if the corresponding weighted filtration $(E_{t\bullet},\alpha_{\bullet})$ satisfies $M(E_{t\bullet},\alpha_{\bullet})=\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=\mu(E_{t\bullet},\alpha_{\bullet},f_{t})=0$.
Proof.
Let $\operatorname{Gies}(t)=(t_{1},t_{2},t_{3})\in\mathbb{G}{(m)}=\mathbb{G}{(m)}_{1}\times\mathbb{G}{(m)}_{2}\times\mathbb{G}{(m)}_{3}$. Recall from Section 6.1 that
$\mu(\operatorname{Gies}(t),\lambda)=s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)+s_{1}\chi\mu(t_{3},\lambda),$
where $s_{1}$ and $s_{2}$ satisfy (28) and $\chi\gg\delta\gg 0$.
First assume that $\mu(\operatorname{Gies}(t),\lambda)=0$. As in the proof of Lemma 4.6, there exists a finite set of weights $\Lambda$ — independent of $t$ — and $g\in\mathrm{GL}(V)$ — depending on $t$ — such that
$\mu(\operatorname{Gies}(t),\lambda)=\sum_{\lambda^{\prime}\in\Lambda}l_{\lambda^{\prime}}\mu(\operatorname{Gies}(t),g\lambda^{\prime}g^{-1}),$
for some $l_{\lambda^{\prime}}\geq 0$. Since $\operatorname{Gies}(t)$ is semistable, $\mu(\operatorname{Gies}(t),g\lambda^{\prime}g^{-1})\geq 0$ for every $\lambda^{\prime}\in\Lambda$. In particular, for each $\lambda^{\prime}\in\Lambda$, either $l_{\lambda^{\prime}}=0$ or
$\mu(\operatorname{Gies}(t),g\lambda^{\prime}g^{-1})=s_{1}\mu(t_{1},g\lambda^{\prime}g^{-1})+s_{2}\mu(t_{2},g\lambda^{\prime}g^{-1})+s_{1}\chi\mu(t_{3},g\lambda^{\prime}g^{-1})=0.$
Since $\chi\gg\delta\gg 0$, i.e. $\chi\gg\epsilon^{-1}\gg 0$, by (28) this implies that $\mu(t_{3},g\lambda^{\prime}g^{-1})=s_{1}\mu(t_{1},g\lambda^{\prime}g^{-1})+s_{2}\mu(t_{2},g\lambda^{\prime}g^{-1})=0$. Therefore, plugging this back into (6.4),
$\mu(t_{3},\lambda)=0=s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda).$
Because of the first equation, the associated filtration $V_{\bullet}$ is $\Gamma$-invariant, and so is the corresponding filtration $E_{t\bullet}$, which implies that $\mu(E_{t\bullet},\alpha_{\bullet},f_{t})=0$.
We claim that $s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)=0$ implies that $M(E_{t\bullet},\alpha_{\bullet})+\delta\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=0$. This is just the same calculation as for the moduli space of principal $G$-bundles with no twisted equivariant structure. For example, use a similar calculation to the one in the proof of [21, Theorem 2.3.5.13] after the statement of [21, Corollary 2.3.5.14].
Next, using Lemma 4.6, we may find a finite set $S$ and a decomposition
$M(E_{t\bullet},\alpha_{\bullet})+\delta\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=\sum_{(\operatorname{rk}E_{\bullet}^{\prime},\operatorname{rk}\alpha_{\bullet}^{\prime})\in S}l_{(E_{\bullet}^{\prime},\alpha_{\bullet}^{\prime})}\left(M(E_{t\bullet}^{\prime},\alpha_{\bullet}^{\prime})+\delta\mu(E_{t\bullet}^{\prime},\alpha_{\bullet}^{\prime},\sigma_{t})\right),$
with $l_{(E_{\bullet}^{\prime},\alpha_{\bullet}^{\prime})}\geq 0$ and only finitely many $l_{(E_{\bullet}^{\prime},\alpha_{\bullet}^{\prime})}>0$. Since $(E_{t},\sigma_{t},f_{t})$ is semistable, either $l_{(E_{\bullet}^{\prime},\alpha_{\bullet}^{\prime})}=0$ or $M(E_{t\bullet}^{\prime},\alpha_{\bullet}^{\prime})+\delta\mu(E_{t\bullet}^{\prime},\alpha_{\bullet}^{\prime},\sigma_{t})=0$. Therefore, for $\delta\gg 0$, either $l_{(E_{\bullet}^{\prime},\alpha_{\bullet}^{\prime})}=0$ or $M(E_{t\bullet}^{\prime},\alpha_{\bullet}^{\prime})=\mu(E_{t\bullet}^{\prime},\alpha_{\bullet}^{\prime},\sigma_{t})=0$. This, in turn, implies that
$M(E_{t\bullet},\alpha_{\bullet})=\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=0,$
as required.
Now assume that $M(E_{t\bullet},\alpha_{\bullet})=\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=\mu(E_{t\bullet},\alpha_{\bullet},f_{t})=0$. In particular $E_{t\bullet}$ is $\Gamma$-invariant, and so is the associated filtration $V_{\bullet}$, which in turn yields $\mu(t_{3},\lambda)=0$. Hence, we have to show that $M(E_{t\bullet},\alpha_{\bullet})+\delta\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=0$ implies $s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)=0$. This is the same calculation as for the moduli space of principal $G$-bundles with no twisted equivariant structure. For example, use a similar calculation to the one in the proof of [21, Theorem 2.3.5.16]. ∎
Proposition 6.10.
A semistable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundle $(Q,\bullet)$ over $X$ is poly-stable if and only if the corresponding point of $\mathfrak{C}_{0}^{ss}$ has closed $\mathrm{GL}(V)$-orbit.
Proof.
Let $t\in\mathfrak{C}_{0}^{ss}$ be the point corresponding to $(Q,\bullet)$, with corresponding pseudo-equivariant pair $(E_{t},\sigma_{t},f_{t})$. For each $t^{\prime}\in\overline{\mathrm{GL}(V)\cdot t}\subset\mathfrak{C}_{0}^{ss}$, denote the corresponding pseudo-equivariant pair by $(E_{t^{\prime}},\sigma_{t^{\prime}},f_{t^{\prime}})$. By Proposition 5.5, we have to show that $(E_{t},\sigma_{t},f_{t})$ is polystable if and only if $(E_{t},\sigma_{t},f_{t})\cong(E_{t^{\prime}},\sigma_{t^{\prime}},f_{t^{\prime}})$ for every $t^{\prime}\in\overline{\mathrm{GL}(V)\cdot t}$.
By the Hilbert–Mumford criterion, there exists a non-trivial 1-parameter subgroup $\lambda:\mathbb{C}^{*}\to\mathrm{GL}(V)$ such that $\lim_{z\to\infty}\lambda(z)\cdot t\in\mathrm{GL}(V)\cdot t^{\prime}$ and $\mu(\operatorname{Gies}(t),\lambda)=0$, where $\operatorname{Gies}$ is the Gieseker map — see Section 6.1. Let $(E_{t\bullet},\alpha_{\bullet})$ be the weighted filtration determined by $\lambda$. By Lemma 6.9, $M(E_{t\bullet},\alpha_{\bullet})=\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=\mu(E_{t\bullet},\alpha_{\bullet},f_{t})=0$. Since $\chi\gg\delta\gg 0$, $\mu(E_{t\bullet},\alpha_{\bullet},f_{t})=\mu(E_{t\bullet},\alpha_{\bullet},\sigma_{t})=M(E_{t\bullet},\alpha_{\bullet})=0$. In particular, $E_{\bullet}$ is $\Gamma$-invariant and the $\mathbb{C}^{*}$-action on $E_{t}$ given by $\lambda$ respects the reduction of $\mathcal{H}om(\mathcal{O}^{\oplus n},E)$ to $G$ given by $\sigma_{t}$.
Therefore, $\lambda$ defines a non-trivial 1-parameter subgroup $\overline{\lambda}$ of a $\Gamma$-invariant maximal torus $T$ of $G$ which, after possibly replacing $t$ with another element in its $\mathrm{GL}(V)$-orbit, we may assume to be independent of $\lambda$. Since $\lambda$ is $\Gamma$-invariant, so is $\overline{\lambda}$. Therefore, $\overline{\lambda}$ is generated by some $s\in i\mathfrak{t}^{\Gamma}$, which in turn defines parabolic subgroup $P_{s}\subset G$ and $\widetilde{P}_{s}\subset\mathrm{GL}(n,\mathbb{C})$. The decomposition of $E$ into weight subspaces with respect to the $\mathbb{C}^{*}$-action given by $\lambda$ defines the weighted filtration $(E_{\bullet},\alpha_{\bullet})$, which in turn determines a reduction of structure group $\widetilde{\tau}\in H^{0}\left(X,\mathcal{H}om(\mathcal{O}^{\oplus n},E)(\mathrm{GL}(n,\mathbb{C})/\widetilde{P}_{s})\right)$ of the frame bundle. This, in turn, determines a reduction $\tau\in H^{0}(X,Q(G/P_{s}))$. Moreover, $\deg E(\tau,s)=M(E_{t\bullet},\alpha_{\bullet})=0$.
It can be seen that $\lim_{z\to\infty}\lambda(z)\cdot E_{t}=\bigoplus_{j}E_{j}/E_{j-1}$, with induced $G$-bundle structure $\lim_{z\to\infty}\lambda(z)\cdot\sigma_{t}$ and equivariant structure $\lim_{z\to\infty}\lambda(z)\cdot f_{t}$. The frame bundle of $\lim_{z\to\infty}\lambda(z)\cdot E_{t}$ is equal to $\mathcal{H}om(\mathcal{O}^{\oplus n},E)_{\widetilde{\tau}}(\widetilde{L}_{s})$, where $\mathcal{H}om(\mathcal{O}^{\oplus n},E)_{\widetilde{\tau}}$ is the $\widetilde{P}_{s}$-bundle over $X$ associated to $\widetilde{\tau}$, and $\widetilde{L}_{s}$ is the Levi subgroup of $\mathrm{GL}(n,\mathbb{C})$ determined by $s$. Therefore, the associated $G$-bundle by $\lim_{z\to\infty}\lambda(z)\cdot\sigma_{t}$ is equal to the extension of structure group of $Q_{\tau}(L_{s})$ to $G$.
We conclude that $(E_{t},\sigma_{t},f_{t})\cong(E_{t^{\prime}},\sigma_{t^{\prime}},f_{t^{\prime}})$ holds if and only if there is a $\Gamma$-equivariant isomorphism $Q_{\tau}(L_{s})\cong Q$, which is true if and only if $Q_{\tau}$ admits a $\Gamma$-invariant reduction of structure group to $L_{s}$ — this last equivalence can be shown using Proposition 5.5. From this it is clear that, if $(Q,\bullet)$ is polystable, then $(E_{t},\sigma_{t},f_{t})\cong(E_{t^{\prime}},\sigma_{t^{\prime}},f_{t^{\prime}})$ for every $t^{\prime}\in\overline{\mathrm{GL}(V)\cdot t}$.
Conversely, assume that $(E_{t},\sigma_{t},f_{t})\cong(E_{t^{\prime}},\sigma_{t^{\prime}},f_{t^{\prime}})$ for every $t^{\prime}\in\overline{\mathrm{GL}(V)\cdot t}$. Take $s\in i\mathfrak{t}^{\Gamma}\setminus\{0\}$ and $\tau\in H^{0}(X,Q(G/P_{s}))^{\Gamma}$ such that $\deg E(\tau,s)=0$ — by Definition 2.3, we have to show that there is a further reduction of structure group of $Q_{\tau}$ to $L_{s}$. The data $\tau,s$ induces a weighted filtration $(E_{t\bullet},\alpha_{\bullet})$. As in Lemma 6.2, we may assume that $h^{1}(X,E_{j}(m))=0$ and $E_{j}(m)$ is globally generated for every $j$, therefore $(E_{t\bullet},\alpha_{\bullet})$ is determined by a non-trivial 1-parameter subgroup $\lambda$ of $\mathrm{GL}(V)$, namely the one associated with the weighted filtration $\left(H^{0}(q_{\mathfrak{C}}\otimes\operatorname{id}_{\pi_{X}^{*}\mathcal{O}_{X}(m)})^{-1}(H^{0}(E_{t\bullet}(m)),\alpha_{\bullet}\right)$ of $V$. Set $t^{\prime}:=\lim_{z\to\infty}\lambda(z)\cdot t$. Then, by the first sentence of the previous paragraph, $(E_{t},\sigma_{t},f_{t})\cong(E_{t^{\prime}},\sigma_{t^{\prime}},f_{t^{\prime}})$ implies that $Q_{\tau}$ admits a reduction of structure group to $L_{s}$, as required. ∎
6.5. The moduli space
Theorem 6.11.
Assume that $G$ is a connected semisimple complex Lie group. Then there exists a complex projective variety $\mathcal{M}(X,G,\Gamma,\theta,c)$ which is a coarse moduli space classifying isomorphism classes of polystable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$. It contains an open subvariety classifying isomorphism classes of stable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$.
Proof.
By Theorem 6.6 and Propositions 4.11 and 3.3, we may choose $m\geq m_{0}$, such that the linearization given in Section 6.1 provides a GIT quotient $\mathcal{M}(X,G,\Gamma,\theta,c):=\mathfrak{C}_{0}^{ss}\sslash\mathrm{GL}(V)$ parametrizing S-equivalence classes of semistable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$. By Proposition 6.10, $\mathcal{M}(X,G,\Gamma,\theta,c)$ classifies isomorphism classes of polystable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$. By Proposition 6.8, it is a complex projective variety. ∎
7. The moduli space for $G$ reductive
Fix the data $X,G,\Gamma,\theta$ and $c$ as in previous sections, where $G$ is now an arbitrary connected reductive complex Lie group. Recall that the construction in Sections 5 and 6 only works when $G$ is semisimple. However, we may use it to build the moduli space of $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles for reductive $G$. We follow the approach in [14], which is a generalization of [22].
We include here a result that will be needed throughout this section. Recall that, in this general setting, the (poly, semi)stability notions depend on a parameter, which is a $\Gamma$-invariant element $\zeta\in i\mathfrak{z}_{\mathfrak{k}}^{\Gamma}$ — here $\mathfrak{z}_{\mathfrak{k}}$ is the centre of a $\Gamma$-invariant maximal compact subalgebra $\mathfrak{k}$ of $\mathfrak{g}$, with corresponding maximal compact subgroup $K$ —, and a $G$-invariant pairing $\mathfrak{g}\cong\mathfrak{g}^{*}$.
Recall that the topology of a $G$-bundle over $X$ can be identified with an element of $Z(G)^{*}$, or $\mathfrak{z}^{*}$, where $Z(G)$ is the centre of $G$ with Lie algebra $\mathfrak{z}$. Via the above pairing, this may be regarded as an element of $\mathfrak{z}$.
Definition 7.1.
The topological type of a $G$-bundle $Q$ is the number $\frac{i}{2\pi}\int_{X}F$, where $F$ is the Chern curvature corresponding to some smooth reduction of structure group of $Q$ to the maximal compact subgroup $K$.
Proposition 7.2.
Take $\zeta\in i\mathfrak{z}_{\mathfrak{k}}^{\Gamma}$ and let $(Q,\bullet)$ be a $\zeta$-polystable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundle over $X$. Then $Q$ has topological type $\zeta$. Conversely, if a $G$-bundle of topological type $\zeta$ is $\zeta_{0}$-polystable for some $\zeta_{0}\in i\mathfrak{z}_{\mathfrak{k}}^{\Gamma}$, then $\zeta_{0}=\zeta$.
Proof.
By Proposition 2.5, $Q$ is $\zeta$-polystable as a $G$-bundle. Therefore, by the Hitchin–Kobayashi correspondence, there exists a smooth reduction of structure group to $K$ whose corresponding Chern curvature $F$ satisfies
$F=-2\pi i\zeta\omega,$
where $\omega$ is a volume form on $X$ with total volume 1. In particular,
$\frac{i}{2\pi}\int_{X}F=\zeta.$
Since the left hand side is the topological type of $Q$, the proposition follows. ∎
7.1. Finite torsors and moduli problems
This section is inspired by [14]. We consider two moduli problems, represented by presheaves $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ over a site of of schemes over $\mathbb{C}$ — with a suitable Grothendieck topology — with values in groupoids. These presheaves represent families of geometric objects over $X$, parametrized by the input scheme. We make the following assumptions.
  1. (1)
    There is a natural transformation $\eta:\mathcal{F}_{1}\to\mathcal{F}_{2}$ and a finite abelian group $A$ such that, for every contractible scheme $U$ — in the strong analytic topology — and every $p\in\mathcal{F}_{2}(U)$, $\eta^{-1}(p)$ is an $A$-torsor.
  2. (2)
    For every complex variety $S$, the functor $\eta(S):\mathcal{F}_{1}(S)\to\mathcal{F}_{2}(S)$ makes $\mathcal{F}_{1}(S)$ a fibred category over $\mathcal{F}_{2}(S)$.
  3. (3)
    For every scheme $S/\mathbb{C}$, $\eta(S):\mathcal{F}_{1}(S)\to\mathcal{F}_{2}(S)$ is surjective on objects.
  4. (4)
    There exist a complex variety $R_{2}$ and an element $u_{2}\in\mathcal{F}_{2}(R_{2})$ such that any family in the image of $\mathcal{F}_{2}$ is locally the pullback of $u_{2}$ by some morphism to $R_{2}$. In other words, the sheafification ~$\widetilde{\mathcal{F}}_{2}$ of $\mathcal{F}_{2}$ is representable by $R_{2}$.
  5. (5)
    There exists a complex reductive group $H$ and an algebraic $H$-action on $R_{2}$ which lifts to an $H$-action on $u_{2}$ and such that, given two points $r$ and $r^{\prime}\in R_{2}$, $u_{2}|_{\{r\}\times X}\cong u_{2}|_{\{r^{\prime}\}\times X}$ if and only if $r^{\prime}\in H\cdot r$.
Lemma 7.3.
Let ~$\widetilde{\mathcal{F}}_{2}$ be the sheafification of $\mathcal{F}_{2}$. Then ~$\widetilde{\mathcal{F}}_{2}$ is representable by a complex variety $R_{1}$ which is an unramified Galois covering $p:R_{1}\to R_{2}$ with Galois group $A$. Moreover, the $H$-action on $R_{2}$ lifts to an $H$-action on $R_{1}$, such that two points $r$ and $r^{\prime}\in R_{1}$ are in the same $H$-orbit if and only if $u_{1}|_{\{r\}\times X}\cong u_{1}|_{\{r^{\prime}\}\times X}$, where $u_{1}\in\widetilde{\mathcal{F}}_{1}(R_{1})$ is a universal family.
Proof.
The construction of $R_{1}$ is similar to [21, Theorem 2.4.8.7]. Given an open subspace $U$ of $R_{2}$, we denote by $\mathcal{F}_{2}^{U}$ the functor from the site of schemes over $U$ to sets, sending a morphism $f:S\to U$ to $\eta(S)^{-1}(f^{*}u_{2}|_{U\times X})$.
Since ~$\widetilde{\mathcal{F}}_{2}$ is a sheaf, it suffices to check that there is an open covering $\curly U=(U_{i})_{i}$ of $R_{2}$ such that $\mathcal{F}_{2}^{U_{i}}$ is representable by $U_{i}\times A$ as an $A$-torsor. This follows from Assumption (2). After gluing, we obtain an unramified Galois cover $p:R_{1}\to R_{2}$ with Galois group $A$. The universal family $u_{1}\in\widetilde{\mathcal{F}}_{1}(R_{1})$ is obtained by gluing elements $u_{1}|_{p^{-1}(U_{i})\times X}\in\mathcal{F}_{1}(p^{-1}(U_{i}))$. There is a natural isomorphism $\widetilde{\eta}(R_{1})(u_{1})\xrightarrow{\sim}p^{*}(u_{2})$, where $\widetilde{\eta}:\widetilde{\mathcal{F}}_{1}\to\widetilde{\mathcal{F}}_{2}$ is induced by $\eta$.
In order to lift the $H$-action, we use the same argument as in the proof of [21, Proposition 2.4.8.9]. The $H$-action on $R_{2}$ is given by a morphism
$\varphi_{H}:H\times R_{2}\to R_{2}.$
The lift of this $H$-action to $u_{2}$, which exists by Assumption (5), is given by an isomorphism
$\widetilde{\varphi}_{H}:\pi_{R_{2}}^{*}u_{2}\xrightarrow{\sim}(\varphi_{H}\times\operatorname{id}_{X})^{*}u_{2}$
over $H\times R_{2}\times X$. Pulling back to $R_{1}$ provides an isomorphism
$\widetilde{\varphi}^{\prime}_{H}:p^{*}\pi_{R_{2}}^{*}u_{2}\xrightarrow{\sim}p^{*}(\varphi_{H}\times\operatorname{id}_{X})^{*}u_{2}$
$H\times R_{1}\times X$. Via the isomorphism $\widetilde{\eta}(R_{1})(u_{1})\xrightarrow{\sim}p^{*}(u_{2})$ and Assumption (2), $\widetilde{\varphi}^{\prime}_{H}$ induces an automorphism of $\pi^{*}_{H}u_{1}$ over $H\times R_{1}\times X$. This, in turn, induces an $H$-action on $R_{1}$.
By construction, if two points of $R_{1}$ are in the same $H$-orbit, then they yield isomorphic objects over $X$. Conversely, if $u_{1}|_{\{r\}\times X}\cong u_{2}|_{\{r^{\prime}\}\times X}$ for two points $r$ and $r^{\prime}\in R_{1}$, then $p^{*}u_{2}|_{\{r\}\times X}\cong p^{*}u_{2}|_{\{r^{\prime}\}\times X}$ and $p(r^{\prime})\in H\cdot p(r)$ by Assumption (5), which by construction implies that $r^{\prime}\in H\cdot r$. ∎
7.2. Moduli space for $G$ abelian
In this section we address the construction of the moduli space of $(\theta,c)$-twisted equivariant $G$-bundles for $G$ abelian. In other words, $G\cong(\mathbb{C}^{*})^{\times p}$ for some $p\in\mathbb{N}$. In this case
(30) $\operatorname{Aut}(G)\cong\mathrm{GL}_{p}(\mathbb{Z}),$
where $\mathrm{GL}_{p}(\mathbb{Z})$ acts on $\mathbb{C}^{*}$ via
$(a_{ij})_{1\leq i,j\leq p}\cdot(z_{1},\dots,z_{p}):=(\prod_{k}z_{k}^{a_{1k}},\prod_{k}z_{k}^{a_{2k}},\dots,\prod_{k}z_{k}^{a_{pk}}).$
Recall that there is a subgroup $Z^{1}_{\theta}(\Gamma,G)\subset G^{\Gamma}$ consisting of maps $\kappa:\Gamma\to G$ which satisfy
$\kappa(\gamma_{1}\gamma_{2})=\kappa(\gamma_{1})\theta_{\gamma_{1}}(\kappa(\gamma_{2})),$
called 1-cocycles. The group $G$ acts on $Z^{1}_{\theta}(\Gamma,G)$ by
$Z^{1}_{\theta}(\Gamma,G)\times G\to Z^{1}_{\theta}(\Gamma,G);\,\kappa\cdot g:=g^{-1}\kappa({\gamma})\theta_{\gamma}(g).$
The quotient $H^{1}_{\theta}(\Gamma,G):=Z^{1}_{\theta}(\Gamma,G)/G$ is the first cohomology group of $\Gamma$ with values in $G$ — see, for example, [23].
Proposition 7.4.
Given a $G$-bundle $Q\to X$, the set of isomorphism classes of $(\theta,c)$-twisted $\Gamma$-equivariant structures on $Q$ is either empty or an $H^{1}_{\theta}(\Gamma,G)$-torsor.
Proof.
Given a $G$-bundle $Q\to X$ admitting some $(\theta,c)$-twisted $\Gamma$-equivariant structure $\bullet$, the set of $(\theta,c)$-twisted $\Gamma$-equivariant structures on $Q$ is a $Z^{1}_{\theta}(\Gamma,G)$-torsor. Indeed, any other $(\theta,c)$-twisted $\Gamma$-equivariant structure * is of the form
$q*\gamma=(q\cdot\kappa(\gamma))\bullet\gamma$
for each $q\in Q$ and $\gamma\in\Gamma$, where $\kappa:\Gamma\to G$ is some map. The condition that * is $(\theta,c)$-twisted — i.e., the corresponding 2-cocycle $c$ is the same as that of $\bullet$ — implies that $\kappa$ is a 1-cocycle in $Z^{1}_{\theta}(\Gamma,G)$.
Moreover, any automorphism of $Q$ is given by multiplication by an element $g\in G$. This fits into the commutative diagramme
${Q}$${Q}$${Q}$${Q}$$\scriptstyle{\cdot g}$$\scriptstyle{(\cdot\kappa({\gamma}))\bullet\gamma}$$\scriptstyle{(\cdot g^{-1}\kappa({\gamma})\theta_{\gamma}(g))\bullet\gamma}$$\scriptstyle{\cdot g}$
for each $\gamma\in\Gamma.$ Therefore, the group of $G$-bundle automorphisms of $Q$ acts on $Z^{1}_{\theta}(\Gamma,G)$ via
$Z^{1}_{\theta}(\Gamma,G)\times G\to Z^{1}_{\theta}(\Gamma,G);\,\kappa\cdot g:=g^{-1}\kappa({\gamma})\theta_{\gamma}(g).$
The set of isomorphism classes of $(\theta,c)$-twisted $\Gamma$-equivariant structures on $Q$ is thus a torsor of the quotient of $Z^{1}_{\theta}(\Gamma,G)$ by this $G$-action, which is precisely equal to the first cohomology group $H^{1}_{\theta}(\Gamma,G)$. ∎
Fix a character $d\in G^{*}$. Via the isomorphism $G^{*}\cong\mathbb{Z}^{p}$, this can be identified with a tuple of $p$ integers $d=(d_{1},\dots,d_{p})$. A $G$-bundle $Q$ has topological type $d$ — according to Definition 7.1 — if and only if it is a product of $p$ $\mathbb{C}^{*}$-bundles of degrees $d_{1},\dots,d_{p}$. Therefore, in the following we call $d$ the degree of $Q$.
The set of isomorphism classes of $G$-bundles of degree $d$ admits a fine moduli space
$\operatorname{Jac}^{d}:=\prod_{1\leq k\leq p}\operatorname{Jac}^{d_{k}},$
where $\operatorname{Jac}^{d_{k}}$ is the degree $d_{k}$ part of the Picard group of $X$. A universal $G$-bundle is given by a product of Poincaré bundles
(31) $\mathcal{P}^{d}:=\bigotimes_{1\leq k\leq p}\pi_{k}^{*}\mathcal{P}^{d_{k}},$
where $\pi_{k}:G\to\mathbb{C}^{*}$ is the projection onto the $k$-th factor. Let $\operatorname{Jac}^{d,\Gamma}_{\theta,c}\subset\operatorname{Jac}^{d}$ be the closed subvariety consisting of $G$-bundles of degree $d$ admitting a $(\theta,c)$-twisted $\Gamma$-equivariant structure. This is a connected component of the subvariety $(\operatorname{Jac}^{d})^{\Gamma}\subset\operatorname{Jac}$ of $\Gamma$-fixed points, where $\gamma\in\Gamma$ sends a $G$-bundle $Q$ to $\gamma^{*}\theta_{\gamma}(Q)$ — here $\theta_{\gamma}(Q)$ is the extension of structure group of $Q$ by $\theta_{\gamma}$. Note that the subvarieties $\operatorname{Jac}^{d,\Gamma}_{\theta,c}$ and $\operatorname{Jac}^{d,\Gamma}_{\theta,c^{\prime}}$ are the same if $c$ and $c^{\prime}$ belong to the same cohomology class in $H^{2}_{\theta}(\Gamma,G)$, and they are disjoint otherwise — see the beginning of Section 2 for the definition of $H^{2}_{\theta}(\Gamma,G)$.
Lemma 7.5.
Fix a complex scheme $S$. The groupoid $\mathcal{F}_{1}(S)$ of families of $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$ parametrized by $S$ is fibred over the groupoid $\mathcal{F}_{2}(S)$ of families of $G$-bundles over $X$ parametrized by $S$ and admitting a $(\theta,c)$-twisted $\Gamma$-equivariant structure. If $S$ is contractible, for every object $L\in\mathcal{F}_{2}(S)$, the set of isomorphism classes of elements in the fibre of $L$ is an $H^{1}_{\theta}(\Gamma,G)$-torsor.
Proof.
The forgetful functor $\mathcal{F}_{1}(S)\to\mathcal{F}_{2}(S)$ makes $\mathcal{F}_{1}(S)$ a fibred category over $\mathcal{F}_{2}(S)$. Indeed, given an isomorphism of $G$-bundles $Q_{1}\to Q_{2}$ and a $(\theta,c)$-twisted $\Gamma$-equivariant structure on $Q_{2}$, the isomorphism induces a $(\theta,c)$-twisted $\Gamma$-equivariant structure on $Q_{1}$ which determines a cartesian diagramme. When $S$ is contractible, the family $\mathcal{F}_{2}(S)$ is isomorphic to $S\times\mathcal{F}_{2}(\operatorname{pt})$ and the fibres are $H^{1}_{\theta}(\Gamma,G)$-torsors by Proposition 7.4. ∎
Theorem 7.6.
Assume that $G\cong(\mathbb{C}^{*})^{p}$. Then the set of isomorphism classes of $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles of degree $d$ over $X$ admits a fine moduli space $\widetilde{\operatorname{Jac}}_{\theta,c}^{d,\Gamma}$, which is an unramified Galois cover of $\operatorname{Jac}^{d,\Gamma}_{\theta,c}\subset\operatorname{Jac}^{d}$ with Galois group $H^{1}_{\theta}(\Gamma,G)$.
Proof.
It is well known that the abelian group $H^{1}_{\theta}(\Gamma,G)$ is finite. Consider the functor $\mathcal{F}_{1}$ sending a complex scheme $S$ to the groupoid of families of $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles of degree $d$ over $X$ parametrized by $S$. Let $\mathcal{F}_{2}$ be the functor sending a complex scheme $S$ to the groupoid of families of $G$-bundles of degree $d$ over $X$ parametrized by $S$, admitting a $(\theta,c)$-twisted $\Gamma$-equivariant structure. By Lemma 7.5, Assumptions (1), (2) and (3) in Section 7.1 are satisfied, by setting $A=H^{1}_{\theta}(\Gamma,G)$. The space $\operatorname{Jac}^{d,\Gamma}_{\theta,c}\subset\operatorname{Jac}^{d}$ is a fine moduli space representing $\mathcal{F}_{2}$, together with a universal family which is the restriction of the Poincaré bundle (31), hence Assumption (4) holds. Assumption (5) is satisfied with $H=1$. Therefore, the theorem follows from Lemma 7.3. ∎
7.3. Moduli space for general reductive $G$
Let now $G$ be an arbitrary connected reductive complex Lie group. Its Lie algebra features a decomposition $\mathfrak{g}=\mathfrak{g}^{\prime}\oplus\mathfrak{z}$, where $\mathfrak{g}^{\prime}:=[\mathfrak{g},\mathfrak{g}]$ and $\mathfrak{z}$ is its centre. Therefore, there exist a semisimple group $G^{\prime}$ — namely, $G^{\prime}=[G,G]$ —, an abelian group $Z$ and an unramified finite Galois covering $p_{G}:G\to G^{\prime}\times Z$, with abelian Galois group $\Lambda:=\operatorname{Gal}(G/G^{\prime}\times Z)$. This restricts to an unramified $\Lambda$-Galois cover $Z(G)\to Z$, where $Z(G)$ is the centre of $G$.
Note that $\theta$ and $c$ induce a homomorphism $\overline{\theta}:\Gamma\to\operatorname{Aut}(G^{\prime}\times Z)$ and a 2-cocycle $\overline{c}\in H^{2}(\Gamma,Z)$.
Since $Z$ is connected, abelian and reductive, $Z\cong(\mathbb{C}^{*})^{\times p}$ for some $p\in\mathbb{N}$. Fix an element $\zeta\in i\mathfrak{z}_{\mathfrak{k}}^{\Gamma}$, where $\mathfrak{z}_{\mathfrak{k}}$ is the centre of a maximal compact subalgebra of $\mathfrak{g}$. By Proposition 7.2, in order to parametrize $\zeta$-polystable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$, we need to focus on $G$-bundles with topological type $\zeta$. Let $d_{k}:=\pi_{k}(\zeta)\in\mathbb{R}$. We actually assume that $d_{k}\in\mathbb{Z}$, since otherwise the moduli space is empty. Then, given a $\zeta$-polystable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundle $(Q,\bullet)$ over $X$, the $\mathbb{C}^{*}$-bundle $\pi_{k}\pi_{Z}p_{G}(Q)$ has degree $d_{k}$, where $\pi_{k}:Z\to\mathbb{C}^{*}$ is the projection onto the $k$-th factor. In other words, using the notation of Section 7.2, $\pi_{Z}p_{G}(Q)$ has degree $d=\zeta$.
Theorem 7.7.
Let $G$ be an arbitrary connected reductive complex Lie group and fix $\zeta\in i\mathfrak{z}_{\mathfrak{k}}^{\Gamma}$. Then there exists a complex projective variety $\mathcal{M}(X,G,\Gamma,\theta,c)$ parametrizing isomorphism classes of $\zeta$-polystable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$. It contains an open subvariety classifying isomorphism classes of stable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$.
Proof.
By [3, Proposition 6.3], the set of isomorphism classes of $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$ is parametrized by a set $H^{1}_{\Gamma,\theta,c}(X,\underline{G})$, which is similar to a Čech cohomology set. Similarly, the set of isomorphism classes of $(\overline{\theta},\overline{c})$-twisted $\Gamma$-equivariant $G^{\prime}\times Z$-bundles over $X$ is parametrized by $H^{1}_{\Gamma,\overline{\theta},\overline{c}}(X,\underline{G^{\prime}\times Z})$. Extension of structure group by $p_{G}$ induces a map
(32) $p_{G}:H^{1}_{\Gamma,\theta,c}(X,\underline{G})\to H^{1}_{\Gamma,\overline{\theta},\overline{c}}(X,\underline{G^{\prime}\times Z}).$
Given an element $(Q,\bullet)\in H^{1}_{\Gamma,\theta,c}(X,\underline{G})$, the fibre $p_{G}^{-1}(p_{G}(Q,\bullet))$ is an $H^{1}_{\Gamma,\theta}(X,\Lambda)$-torsor by [3, Proposition 6.7], where $H^{1}_{\Gamma,\theta}(X,\Lambda)$ is finite because $\Lambda$ is finite.
Given a $(\overline{\theta},\overline{c})$-twisted $\Gamma$-equivariant $G^{\prime}\times Z$-bundle $(\overline{Q},\bullet)\in H^{1}_{\Gamma,\overline{\theta},\overline{c}}(X,\underline{G^{\prime}\times Z})$, the underlying $G^{\prime}\times Z$-bundle $\overline{Q}$ determines an element $\delta(\overline{Q})\in H^{2}(X,\Lambda)$, such that $\delta(\overline{Q})=0$ if and only if there exists a $G$-bundle $Q$ such that $p_{G}(Q)\cong\overline{Q}$. Moreover, if $\delta(\overline{Q})=0$, the twisted equivariant action $\bullet$ on $Q$ lifts to a $(\theta,c^{\prime})$-twisted $\Gamma$-equivariant action on $Q$, where $[c^{\prime}]\in H^{2}_{\theta}(\Gamma,Z(G))$ is a cohomology class such that $p_{G}(c^{\prime})=\overline{c}$. Multiplying by $c^{-1}$, we get a class $[c^{-1}c^{\prime}]\in H^{2}_{\theta}(\Gamma,Z(G))$. Since $c^{\prime}$ and $c$ induce the same cohomology class $[\overline{c}]\in H^{2}_{\overline{\theta}}(\Gamma,Z)$, the class $[c^{-1}c^{\prime}]$ determines an element of $H^{2}_{\overline{\theta}}(\Gamma,\Lambda)$. Introducing the second cohomology class $H^{2}_{\Gamma,\theta}(X,\Lambda)$ given by [3, Definition 6.6], we therefore find a map
$\delta_{c}:H^{1}_{\Gamma,\overline{\theta},\overline{c}}(X,\underline{G^{\prime}\times Z})\to H^{2}_{\Gamma,\theta}(X,\Lambda)$
containing the data of $\delta(\overline{Q})\in H^{2}(X,A)$ and the class $[c^{-1}c^{\prime}]$. This satisfies that $(\overline{Q},\bullet)\in H^{1}_{\Gamma,\overline{\theta},\overline{c}}(X,\underline{G^{\prime}\times Z})$ is in the image of (32) if and only if $\delta_{c}(\overline{Q},\bullet)=0$.
Consider the functor $\mathcal{F}_{1}$ sending a complex scheme $S$ to the groupoid of families of $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles of topological type $\zeta$ over $X$ parametrized by $S$. Let $\mathcal{F}_{2}$ be the functor sending a complex scheme $S$ to the groupoid of families of $(\overline{\theta},\overline{c})$-twisted $\Gamma$-equivariant $G^{\prime}\times Z$-bundles of topological type $\zeta$ over $X$ parametrized by $S$, whose image by $\delta_{c}$ is equal to 00.
Let us check that Assumptions (1)-(5) in Section 7.1 are satisfied.
  1. (1)
    Extension of structure group by $p_{G}$ determines a natural transformation $\eta:\mathcal{F}_{1}\to\mathcal{F}_{2}$, whose fibres are $H^{1}_{\Gamma,\theta}(X,\Lambda)$-torsors. Therefore, Assumption (1) holds with $A=H^{1}_{\Gamma,\theta}(X,\Lambda)$.
  2. (2)
    Choose a complex scheme $S$, two isomorphic families of $(\overline{\theta},\overline{c})$-twisted $\Gamma$-equivariant $G^{\prime}\times Z$-bundles $(\overline{Q}_{1},\bullet)$ and $(\overline{Q}_{2},\bullet)$ on $S\times X$, a $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundle $(Q_{2},\bullet)$ on $S\times X$ such that $p_{G}(Q_{2},\bullet)=(\overline{Q}_{2},\bullet)$ and an isomorphism
    (33) $(\overline{Q}_{1},\bullet)\xrightarrow{\sim}(\overline{Q}_{2},\bullet)$
    Since $\delta_{G}(\overline{Q}_{1})=0$, there exists a $G$-bundle $Q_{1}$ such that $p_{G}(Q_{1})=\overline{Q}_{1}$. The total spaces of $Q_{1}$ and $Q_{2}$ are finite unramified Galois covers of the total spaces of $\overline{Q}_{1}$ and $\overline{Q}_{2}$, respectively, with the same Galois group $\Lambda$. Therefore, the isomorphism (33) lifts to an isomorphism $Q_{1}\xrightarrow{\sim}Q_{2}$. If we endow $Q_{1}$ with the twisted equivariant structure induced by the twisted equivariant structure on $Q_{2}$ via this isomorphism, then we find a lift $(Q_{1},\bullet)$ of $(\overline{Q}_{1},\bullet)$ completing the cartesian diagramme. Therefore, $\mathcal{F}_{1}(S)$ is a fibred category over $\mathcal{F}_{2}(S)$, so Assumption (2) holds.
  3. (3)
    The functor $\eta(S):\mathcal{F}_{1}(S)\to\mathcal{F}_{2}(S)$ is surjective on objects, since by construction all the objects in $\mathcal{F}_{2}(S)$ have trivial image under $\delta_{c}$.
  4. (4)
    Let $\mathfrak{C}_{0}^{ss}$ be the complex variety satisfying Corollary 6.7. Let $\widetilde{\operatorname{Jac}}_{\overline{\theta},\overline{c}}^{\zeta,\Gamma}$ be the moduli space of $(\theta,c)$-twisted $\Gamma$-equivariant $Z$-bundles of degree $\zeta$ over $X$, given by Theorem 7.6. The map $\delta_{c}$ induces an algebraic morphism
    $\delta_{c}:\mathfrak{C}_{0}^{ss}\times\widetilde{\operatorname{Jac}}_{\overline{\theta},\overline{c}}^{\zeta,\Gamma}\to H^{2}_{\Gamma,\theta}(X,\Lambda).$
    Let $\delta_{c}^{-1}(0)\subset\mathfrak{C}_{0}^{ss}\times\widetilde{\operatorname{Jac}}_{\overline{\theta},\overline{c}}^{\zeta,\Gamma}$. Since $H^{2}_{\Gamma,\theta}(X,\Lambda)$ is finite, this is a union of connected components of $\mathfrak{C}_{0}^{ss}\times\widetilde{\operatorname{Jac}}_{\overline{\theta},\overline{c}}^{\zeta,\Gamma}$. It comes equipped with a universal family $(\mathcal{Q}\times\mathcal{P}^{\zeta},\bullet)$. Set $H=\mathrm{GL}(V)$. By Corollary 6.7, the complex variety $\delta_{c}^{-1}(0)$, together with the $\mathrm{GL}(V)$-action on $\delta^{-1}(0)$ induced by the action on the first factor $\mathfrak{C}_{0}^{ss}$ given by Corollary 6.7, satisfies Assumptions (4) and (5).
By Lemma 7.3, there exists an unramified Galois covering $R_{1}\to R_{2}$, together with a locally universal family over $R_{1}$ and a linearization of the $\mathrm{GL}(V)$-action. Using that $\Lambda$ is finite, it is straightforward to check that a $(\theta,c)$-twisted equivariant $G$-bundle is $\zeta$-(poly, semi)-stable if and only if the corresponding $(\overline{\theta},\overline{c})$-twisted equivariant $G^{\prime}\times Z$-bundle is $\zeta$-(poly, semi)-stable. Since $\mathcal{M}(X,G^{\prime}\times Z,\Gamma,\theta,c):=R_{2}\sslash\mathrm{GL}(V)$ is a coarse moduli space parametrizing isomorphism classes of $\zeta$-polystable $(\overline{\theta},\overline{c})$-twisted $\Gamma$-equivariant $G^{\prime}\times Z$-bundles over $X$, the GIT quotient $\mathcal{M}(X,G,\Gamma,\theta,c):=R_{1}\sslash\mathrm{GL}(V)$ is a coarse moduli space parametrizing isomorphism classes of $\zeta$-polystable $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles over $X$. Moreover, $\mathcal{M}(X,G,\Gamma,\theta,c)$ is a Galois cover of $\mathcal{M}(X,G^{\prime}\times Z,\Gamma,\theta,c)$ with Galois group $H^{1}_{\Gamma,\theta}(X,\Lambda)$. Since $\mathcal{M}(X,G^{\prime}\times Z,\Gamma,\theta,c)$ is projective, so is $\mathcal{M}(X,G,\Gamma,\theta,c)$. ∎